### v2

# Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models

Abstract
Assessing the effectiveness of large language models (LLMs) in addressing diverse tasks is essential for comprehending their strengths and weaknesses. Conventional evaluation techniques typically apply a single prompting strategy uniformly across datasets, not considering the varying degrees of task complexity. We introduce the Hierarchical Prompting Taxonomy (HPT), a taxonomy that employs a Hierarchical Prompt Framework (HPF) composed of five unique prompting strategies, arranged from the simplest to the most complex, to assess LLMs more precisely and to offer a clearer perspective. This taxonomy assigns a score, called the Hierarchical Prompting Score (HP-Score), to datasets as well as LLMs based on the rules of the taxonomy, providing a nuanced understanding of their ability to solve diverse tasks and offering a universal measure of task complexity. Additionally, we introduce the Adaptive Hierarchical Prompt framework, which automates the selection of appropriate prompting strategies for each task. This study compares manual and adaptive hierarchical prompt frameworks using four instruction-tuned LLMs, namely Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B, across four datasets: BoolQ, CommonSenseQA (CSQA), IWSLT-2017 en-fr (IWSLT), and SamSum. Experiments demonstrate the effectiveness of HPT, providing a reliable way to compare different tasks and LLM capabilities. This paper leads to the development of a universal evaluation metric that can be used to evaluate both the complexity of the datasets and the capabilities of LLMs. The implementation of both manual HPF and adaptive HPF is publicly available [^1] GitHub repository consisting of the code implementation of this work: https://github.com/devichand579/HPT.
Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models
Devichand Budagam1, Sankalp KJ2, Ashutosh Kumar3, Vinija Jain4, 5thanks: Work does not relate to position at Amazon, Aman Chadha4, 5thanks: Work does not relate to position at Amazon 1Indian Institute of Technology Kharagpur, India, 2AI Institute, University of South Carolina, USA 3Rochester Institute of Technology, USA, 4Amazon GenAI, USA, 5Stanford University, USA
1 Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements across a wide range of applications. However, evaluating the capabilities of LLMs across various datasets remains a challenging task. Traditional prompting strategies, which apply the same prompt to all dataset samples regardless of task complexity, often result in sub-optimal evaluations. This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of task complexity.
Figure 1: Hierarchical Prompt Framework (HPF) consists of five distinct prompting strategies, each tailored to different levels of task complexity, ensuring the appropriate prompt is selected for the given task.
Prompting plays a vital role in unlocking the full potential of LLMs. By designing specific input prompts, the LLM’s responses can be guided, significantly influencing the quality and relevance of the output. Effective prompting strategies have enhanced LLM performance on tasks ranging from simple question-answering to complex reasoning and problem-solving. Recent research has explored various approaches to prompting and reasoning evaluation in LLMs. Chain-of-Thought (CoT) prompting Wei et al. (2022b) elicits step-by-step reasoning, improving performance on complex tasks. Specializing smaller models Fu et al. (2023) and using large models as reasoning teachers Ho et al. (2022) have demonstrated the potential for enhancing reasoning capabilities. Progressive-hint prompting Zheng et al. (2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses. Metacognitive prompting Wang and Zhao (2024) incorporates self-aware evaluations to enhance understanding abilities.
We introduce the Hierarchical Prompting Taxonomy (HPT), a set of rules that employs a Hierarchical Prompt Framework (HPF) shown in Figure 1, a structured evaluation method addressing the need for more sophisticated assessment of LLM task-solving capabilities. This framework employs five distinct prompting strategies, each tailored to different levels of task complexity. By assigning a score called the HP-Score based on the LLM’s performance across these levels, this work provides a comprehensive assessment of its ability to solve diverse tasks.
Furthermore, the taxonomy incorporates an Adaptive Hierarchical Prompt framework, enabling an LLM, called prompt-selector to dynamically select the appropriate prompting strategy based on a given task complexity. This adaptive approach automates the evaluation process, also provides a better understanding of the judgment skills of LLMs in determining task complexity. The main contributions of this paper are as follows:
  1. 1.
    We propose the Hierarchical Prompting Taxonomy (HPT), a set of rules designed to establish a universal measure of task complexity for both datasets and LLMs.
  2. 2.
    Based on the rules of the HPT, this work introduces a set of five unique prompting strategies, arranged from simplest to most complex, to provide a more precise assessment of LLMs and offer clearer insights.
  3. 3.
    We introduce the Adaptive HPF that allows us to examine the judgement skills of LLMs by automating the evaluation process and dynamically selecting the right prompting approach based on task complexity.
The rest of the paper is organized as follows: Section 2 provides an overview of related work on prompting and evaluation in LLMs. Section 3 presents HPT and the frameworks associated with it in detail. Section 4 describes the experimental setup and results. Section 5 concludes the paper, Section 6 discusses the limitations and finally, Section 7 discusses the ethical impact of this work.
2 Related Work
The advent of LLMs has revolutionized NLP by demonstrating significant improvements in few-shot and zero-shot learning capabilities. Brown et al. (2020) introduced GPT-3, a 175 billion parameter autoregressive model, showcasing its ability to perform a wide range of tasks such as question-answering, reading comprehension, translation, and natural language inference without fine-tuning. This study highlighted the potential of very large models for in-context learning, while also identifying limitations in commonsense reasoning and specific comprehension tasks. Similarly, Liu et al. (2021) surveyed prompt-based learning, emphasizing the role of prompt engineering in leveraging pre-trained models for few-shot and zero-shot adaptation to new tasks with minimal labeled data.
Emergent abilities in LLMs, which appear suddenly at certain scale thresholds, have also been a topic of interest. Wei et al. (2022a) examined these abilities in few-shot prompting, discussing the underlying factors and implications for future scaling. Complementing this, Kojima et al. (2022) demonstrated that LLMs could exhibit multi-step reasoning capabilities in a zero-shot setting by simply modifying the prompt structure, thus highlighting their potential as general reasoning engines. Further exploring reasoning capabilities, Qiao et al. (2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning.
Several studies have focused on improving the interaction and decision-making abilities of LLMs through innovative prompting techniques. Lo et al. (2023) proposed ASH (Actor-Summarizer-Hierarchical) prompting, which enhances LLM performance on complex web navigation tasks by decomposing the process into summarization and action prediction. This hierarchical approach significantly outperformed previous methods on tasks with lengthy trajectories. Kong et al. (2024) introduced role-play prompting to improve zero-shot reasoning by constructing role-immersion interactions, which implicitly trigger chain-of-thought processes and enhance performance across diverse reasoning benchmarks.
The robustness and safety of LLM outputs have also been addressed in recent research. Wang et al. (2023) developed INDUST, a benchmark to evaluate LLM resistance to inductive instructions containing false premises, showing that self-critique prompting can help models handle such challenging instructions more effectively. Additionally, Chen et al. (2024) provided a comprehensive survey on prompt engineering techniques aimed at reducing hallucinations and enhancing the reliability of LLM outputs through advanced methods like chain-of-thought and tree-of-thought prompting.
To facilitate the evaluation and understanding of LLM capabilities, Zhu et al. (2024) introduced PromptBench, a unified library encompassing a variety of LLMs, datasets, evaluation protocols, and adversarial prompt attacks. This modular and extensible tool aims to support collaborative research and advance the comprehension of LLM strengths and weaknesses. Zhou et al. (2023b) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
Chang et al. (2022) proposed SpeechPrompt, a framework that reformulates tasks as conditional generation problems for pre-trained generative spoken language models. This method demonstrated competitive performance in speech classification with fewer trainable parameters and potential in more challenging tasks like automatic speech recognition.
Lastly, Yao et al. (2023) introduced the Tree-of-Thoughts framework, enabling LLMs to deliberate over coherent text units and perform heuristic searches for complex reasoning tasks. This approach generalizes over chain-of-thought prompting and has shown significant performance improvements in tasks requiring planning and search, such as creative writing and problem-solving games.
These works collectively highlight the advancements in leveraging LLMs through innovative prompting techniques, addressing their emergent abilities, reasoning capabilities, interaction strategies, robustness, and evaluation methodologies. Despite significant advancements, the current research on LLMs reveals several limitations, particularly in terms of prompt design, handling complex reasoning tasks, and evaluating model performance across diverse scenarios. The emergent abilities of LLMs, while promising, often lack predictability and control, and the robustness of these LLMs in the face of misleading prompts remains a concern.
To address these challenges, we propose the HPT and HPF for evaluating different LLMs. This approach aims to systematically categorize and assess the effectiveness of various prompting techniques based on task complexity, enabling a more nuanced understanding of LLM behaviors and performance. By introducing a structured framework for evaluation, this work seeks to enhance the reliability and applicability of LLMs, facilitating their deployment in increasingly complex and dynamic environments.
3 Methodology
Figure 2: Hierarchical Prompting Taxonomy (HPT): A taxonomy designed to assess task complexity based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
3.1 Datasets and Models
Experiments were performed on BoolQ Clark et al. (2019), CommonSenseQA (CSQA) Talmor et al. (2019), IWSLT-2017 en-fr (IWSLT) Cettolo et al. (2017), and SamSum Gliwa et al. (2019). BoolQ consists of approximately 16,000 True/False questions with corresponding passages from Wikipedia, designed for binary classification in reading comprehension. CSQA contains around 12,000 multiple-choice questions sourced from ConceptNet, aimed at evaluating model’s commonsense reasoning abilities. IWSLT is a parallel corpus with tens of thousands of English-French sentence pairs from TED Talks, used for machine translation tasks. SamSum features around 16,000 human-generated chat logs with summaries, facilitating dialogue summarization model development. On the aforementioned datasets, the performance of the instruction-tuned LLMs Llama 3 8B AI@Meta (2024), Phi 3 3.8B Abdin et al. (2024), Mistral 7B Jiang et al. (2023), and Gemma 7B Team et al. (2024) were evaluated on their validation splits, with the exception of SamSum, which utilized the test split and nearly 2300 hours of computation was used for all experiments on the V100 GPU.
3.2 Hierarchical Prompting Taxonomy (HPT)
The proposed HPT shown in Figure 2 is founded on principles that are crucial for humans to address the majority of reasoning and interpretation tasks, taking into account the following criteria:
  • Basic Recall and Reproduction: Remember and reproduce factual information without interpretation or analysis.
  • Understanding and Interpretation: Comprehend and explain the meaning of information, summarizing or clarifying content.
  • Analysis and Reasoning: Break down complex information, understand relationships, and solve problems using logical reasoning.
  • Application of Knowledge and Execution: Apply knowledge in practical situations, execute multi-step processes, and solve complex tasks.
3.3 Hierarchical Prompt Framework (HPF)
The HPT employs an HPF that comprises five different prompting strategies. The complexity levels of these prompting strategies within the framework are determined by the taxonomy criteria, which judges actions such as the requirement for external knowledge, the capacity to solve subproblems, and the capacity to provide step-by-step reasoning that takes the task’s complexity into account. The five different prompting strategies that make up the HPF are :
  • Role Prompting Kong et al. (2024): Prompts that define a role for the LLM in solving the task.
  • Zero-Shot Chain-of-Thought Prompting (Zero-CoT) Kojima et al. (2022): Prompts that use the phrase "Let’s think step by step" without providing prior examples to encourage critical thinking and problem-solving.
  • Three-Shot Chain-of-Thought Prompting (3-CoT) Wei et al. (2022b): Prompts that offer three examples to guide the LLM’s reasoning process.
  • Least-to-Most Prompting Zhou et al. (2023a): Prompts that sequentially breakdown the task into sub problems to derive essential insights from the task in order to solve it.
  • Generated Knowledge Prompting (GKP) Liu et al. (2022): Prompts that require the integration of external knowledge to accomplish the task. Llama 3 8B was used to generate the necessary knowledge. Unlike the original study, the knowledge prediction with the highest confidence score was selected among three possible predictions.
The Manual HPF systematically iterates through the predefined levels of prompting strategies to solve a given task using the language models. This process involves starting with the simplest prompts and gradually progressing to more complex ones based on the prompt hierarchy levels until the task is successfully addressed. By manually traversing each level, it ensures that every potential strategy is explored and allowing the LLM to identify the most effective approach for the task.
A prompt-selector determines which prompting strategy is most suited for a given task in the adaptive HPF. This allows us to examine how LLMs may be utilised to assess their abilities to do different tasks while also automating the evaluation process.
3.4 Hierarchical Prompting Score (HP-Score)
The HP-Score is an evaluation metric for assessing the complexity of a task in comparison to various categories of problem-solving agents, such as humans and LLMs, which is heavily influenced by HPT rules. A task’s lower HP-Score indicates that the corresponding agent is more capable of solving the task.
3.4.1 HP-Score for Datasets
Datasets are evaluated by scoring each sample of a representative set of 5% of the original evaluation dataset on a scale from 1 (lowest) to the number of prompting levels (highest) existing in HPF for each of the four HPT rules. HP-Score for datasets is computed as:
$\text{hp}_{\text{j}}=\frac{1}{4}\sum_{i=1}^{4}a_{i}$ (1)
[…]
where $a_{i}$ represents the individual score for each rule of HPT, $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$ and $n$ denotes the total number of samples in the representative set. The HP-Scores assigned by experts in the domains related to the datasets used in the experiments are displayed in Table 1. CSQA is the most complex task relative to humans, with the highest HP-Score among the four datasets.
Dataset Evaluation Set Size Representative Set Size HP-Score
BoolQ 3270 162 1.71
CSQA 1221 60 2.52
IWSLT 890 45 1.92
SamSum 819 40 2.23
Table 1: HP-Scores of datasets by human experts.
3.4.2 HP-Score for Manual HPF
HPF categorizes prompting strategies based on their complexity into five levels: Role Prompting (level 1), Zero-CoT (level 2), 3-CoT (level 3), Least-to-Most Prompting (level 4), and GKP (level 5). Each level has a score of the same value as the level, i.e. for level x, score = x. A LLM’s superior task-solving ability is indicated by a lower HP-Score. If the LLM fails to address the task at all levels of prompts, it is assigned a penalty corresponding to its dataset HP-Score. If the LLM was able to address the task at level x:
$\text{hp}_{\text{j}}=x$ (3)
If LLM was unable to address the task after all m levels:
$\text{hp}_{\text{j}}=m+\text{HP-Score}_{\text{Dataset}}$ (4)
HP-Score for Manual HPF is:
[…]
where $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$, x represents the level of HPF at which task is addressed, $m$ represents the total number of levels in HPF, and $n$ denotes the total number of samples in the evaluation set.
3.4.3 HP-Score for Adaptive HPF
The prompt-selector can dynamically select the most suitable prompting strategy for a given task’s complexity from the manual HPF’s collection of prompting strategies. In order to determine the most effective prompting strategy to address the task, the prompt-selector was given a maximum number of iterations equivalent to the number of levels in the manual HPF. The score for ith iteration is i + x, where x is the level of the prompting strategy selected by the prompt-selector. If the LLM fails to address the task after all iterations, it is assigned a penalty corresponding to it’s dataset HP-Score. If LLM was able to address the task at ith iteration:
$\text{hp}_{\text{j}}=x+i$ (6)
If LLM was unable to address the task after all m iterations:
$\text{hp}_{\text{j}}=m+\text{HP-Score}_{\text{Dataset}}$ (7)
HP-Score for Adaptive HPF is:
[…]
where $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$, x represents the level of HPF selected by prompt-selector at $i$th iteration at which task is addressed, $m$ represents the total number of levels in HPF, and $n$ denotes the total number of samples in the evaluation set.
4 Results
4.1 Results for Manual HPF
Table 2 presents the HP-Scores for the four LLMs: Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B, across previously mentioned datasets with varying threshold values (0.15 and 0.20) for BLEU Papineni et al. (2002) and ROUGE Lin (2004). The BLEU and ROUGE thresholds at every prompting level serve as the task-solving criteria for each sample of IWSLT and SamSum respectively. The complexity at which the LLMs are able to handle the dataset effectively is reflected in the HP-Scores.
Figure 3: HP-Scores of all datasets and threshold values for LLMs in Manual HPF.
Model BoolQ CSQA IWSLT (0.15) IWSLT (0.20) SamSum (0.15) SamSum (0.20)
Llama 3 8B 1.3692 2.4281 3.3955 3.9246 1.3018 1.7166
Mistral 7B 1.4063 2.4929 3.5183 4.1455 1.2590 1.6770
Phi 3 3.8B 1.3965 2.1714 3.8479 4.4401 2.0925 3.4010
Gemma 7B 1.4558 2.5053 5.3929 5.8369 3.3137 5.0316
Table 2: HP-Scores of datasets (with thresholds) across all LLMs for Manual HPF.
Dataset Metric Threshold Llama 3 8B Phi 3 3.8B Mistral 7B Gemma 7B
BoolQ Accuracy - 0.99296 0.98685 0.98073 0.99419
CSQA Accuracy - 0.84766 0.88452 0.82063 0.83783
IWSLT BLEU 0.15 0.23539 0.19327 0.21629 0.08049
0.2 0.23091 0.19098 0.21569 0.06807
SamSum ROUGE-1 0.15 0.29395 0.21389 0.28529 0.16067
0.2 0.29996 0.22109 0.29365 0.14063
Table 3: Evaluation scores for datasets across all LLMs for Manual HPF.
Figure 3 shows that Llama 3 8B achieves lower HP-Scores than other LLMs across all datasets, indicating superior performance and the ability to solve tasks with simpler prompting strategies. All four LLMs were better at solving BoolQ than humans, with lower HP-Scores than BoolQ HP-Score by human experts. IWSLT was the task where all four LLMs were incapable of solving even at threshold values of 0.15 and 0.20, indicating that LLMs are incapable in solving domain-specific tasks compared to human experts in the respective domains.
Table 3 shows the evaluation scores for each LLM across all datasets using the Manual HPF. Llama 3 8B and Gemma 7B outperform Phi 3 3.8B and Mistral 7B on BoolQ, with all four LLMs achieving near-perfect accuracy. Phi 3 3.8B outperformed all other LLMs in solving CSQA. Despite having the fewest parameters, phi 3 3.8B achieved the lowest HP-Score among all four LLMs while also achieving the highest accuracy. Llama 3 8B has the highest BLEU scores and ROUGE scores across both threshold values for IWSLT and SamSum datasets respectively. While Gemma 7B has the lowest performance and the highest HP-Scores for the IWSLT and SamSum datasets. Llama 3 8B and Mistral 7B are the most capable and genearlized LLMs, achieving superior performance across all datasets. Phi 3 3.8B, even with the fewest parameters, was able to compete with Llama 3 8B and Mistral 7B. Gemma 7B even though was able to solve BoolQ and CSQA with superior performance showing it’s reasoning abilities but was unable to achieve good performance across domain-specific datasets like IWSLT and SamSum.
Model BoolQ CSQA IWSLT (0.15) IWSLT (0.20) SamSum (0.15) SamSum (0.20)
Llama3 5.2173 5.9136 6.2006 6.2841 5.0316 5.5756
Mistra 7B 5.0483 5.9073 6.2478 6.4604 4.7423 5.1336
Phi 3 3.8B 5.1386 5.6793 6.3955 6.4936 5.0961 5.7778
Gemma 7B 5.1514 5.5771 6.5947 6.6605 5.7229 6.4347
Table 4: HP-Scores of datasets (with thresholds) across all LLMs for Adaptive HPF.
Dataset Metric Threshold Llama3 8B Phi 3 3.8B Mistral 7B Gemma 7B
BoolQ Accuracy - 0.88577 0.91115 0.91752 0.91166
CSQA Accuracy - 0.59451 0.68019 0.60111 0.68549
IWSLT BLEU 0.15 0.21140 0.15557 0.20000 0.08447
0.2 0.21146 0.15354 0.20568 0.07730
SamSum ROUGE-1 0.15 0.24407 0.20586 0.26910 0.16023
0.2 0.24981 0.21580 0.28335 0.16179
Table 5: Evaluation scores for datasets across all LLMs for Adaptive HPF.
4.2 Results for Adaptive HPF
The Adaptive HPF automates the selection of appropriate prompting strategy for each task using a prompt-selector utilizing Llama 3 8B for the experiments. Table 4 shows the HP-Scores and Table 5 presents the evaluation scores for the LLMs using the adaptive HPF.
Figure 4: HP-Scores of all datasets and threshold values for LLMs in Adaptive HPF.
The findings presented in Figure 4 show that, on all datasets, the adaptive HPF produces HP-Scores for the LLMs that are higher than those obtained with the manual HPF, with decreased evaluation scores. Higher HP-Scores demonstrate that adaptive HPF is obviously unable to complete a variety of tasks, with the primary cause being hallucinations when choosing the appropriate level of prompting by the prompt-selector. Hallucination of prompt-selector leads to repeatedly choosing irrelevant prompting levels or choosing levels out of the range of manual HPF for several iterations of adaptive HPF, resulting in high HP-Scores. This demonstrates that LLMs are incapable of judging their evaluations of various tasks.
Figure 5: Accuracy scores of BoolQ and CSQA datasets across all LLMs for manual HPF and Adaptive HPF.
Figure 6: BLEU scores of IWSLT dataset across all LLMs for manual HPF and Adaptive HPF..
Figure 7: ROUGE-1 scores of SamSum dataset across all LLMs for manual HPF and Adaptive HPF..
4.3 Comparison between Manual HPF and Adaptive HPF
Figure 5 illustrates that the manual HPF outperforms the adaptive HPF on the BoolQ dataset across all LLMs. The decrease in accuracy with the adaptive HPF demonstrates its inability to select the most appropriate prompting level within its iteration limit. Figure 5 shows a significant decline in performance for LLMs in solving CSQA, suggesting that manual HPF is more dependable than adaptive HPF for solving CSQA. Despite solving the IWSLT with slightly worse performance as shown in Figure 6 when using the adaptive HPF, LLMs are unable to identify the optimal prompting level indicated by their high HP-Scores. Figure 7 shows that both frameworks perform similarly in solving SamSum, with the adaptive HPF showing higher HP-Scores across all LLMs. Table 1 shows that the HP-Scores of both manual HPF and adaptive HPF indicate that manual HPF more closely matches human estimation of task complexity, while the adaptive HPF faces difficulty in solving these tasks when compared to both manual HPF and humans. Both frameworks have higher HP-Scores when solving IWSLT than domain experts’ HP-Scores, indicating that pre-trained LLMs with more domain-specific knowledge are needed to demonstrate better performance across domain-specific tasks with less complex prompting strategies, reducing the need for these LLMs to be fine-tuned. Another important finding from the IWSLT and SamSum evaluation scores is that all four LLMs were unable to perform above a particular evaluation score at higher threshold values, as evidenced by the saturation of their performance across all threshold values.
Overall, our experiments demonstrate that different tasks don’t always require sophisticated prompting strategies from LLMs; it largely depends on the LLM’s aptitude for the task. The HP-Score provides a universal measure of task complexity relative to different agents, offering insights into the capabilities of both LLMs and humans. When comparing Manual HPF and Adaptive HPF, the Manual HPF is more reliable and consistently performs well on datasets with lower HP-Scores. In contrast, the Adaptive HPF is computationally inefficient due to hallucinations when selecting the prompting level in least number of iterations. For other tasks and LLMs, Adaptive HPF requires finetuning of the prompt selector or other techniques to minimize hallucinations and ensure consistent outcomes.
5 Conclusion
The HPT, combined with the HPFs, offers a robust and efficient method for evaluating LLMs based entirely on task complexity. These findings highlight the effectiveness of the HPF in evaluating the task with the most effective prompting strategies based on task complexity, leading to enhanced LLM performance across various datasets and tasks. This approach not only provides a detailed understanding of LLMs problem-solving capabilities but also introduces Adaptive HPF by dynamically selecting the most appropriate prompting strategy, paving the way for future advancements in LLM evaluation techniques.
6 Limitations
Restricted Model Evaluation: The study focused on four specific instruction-tuned LLMs: Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B. While these LLMs provide valuable insights, the inclusion of a more diverse set of LLMs, both in terms of architecture and scale, would help assess the applicability of the HPT across a broader spectrum of LLMs.
Restricted Dataset Evaluation: The evaluation of both HPF and adaptive HPF is conducted on four distinct datasets encompassing tasks such as question-answering, reading comprehension, translation, and summarization. Incorporating a wider variety of datasets that cover additional tasks would aid in evaluating the HPT’s applicability across a more extensive range of tasks.
Manual framework optimization: The effectiveness of the HPF heavily relies on the quality of the prompts used at each level of the taxonomy. Crafting high-quality prompts that accurately reflect the subtleties of each level demands considerable expertise and repeated refinement. This study only investigated a limited set of prompting strategies within the HPF, indicating a need for further research into creating diverse structural frameworks and incorporating additional prompting strategies.
Adaptive framework optimization: The Adaptive HPF automates the selection of suitable prompting strategies according to task complexity. However, the current system depends on a prompt-selector that utilizes an LLM. There were numerous hallucinations when choosing the appropriate prompting strategy for the task. Further investigation is required to minimize hallucinations in the prompt-selector and to explore alternative methods, such as ensemble techniques or meta-learning approaches, to enhance the efficiency and effectiveness of the adaptive framework.
7 Ethical Consideration
The HP-Scores assigned by experts in the domains related to the datasets—BoolQ, CSQA, IWSLT and SamSum—may introduce bias into the comparative analysis. This potential bias stems from the subjective nature of expert scoring, which can be influenced by individual experience and perspective. Despite this, the datasets utilized in this study are publicly available and widely recognized in the research community, thereby minimizing the risk of unanticipated ethical issues. Nevertheless, it is crucial to acknowledge the possibility of scoring bias to ensure transparency and integrity in the analysis.
Appendix A Large Language Models Used for Evaluation
The HPF supports multiple LLMs and includes mechanisms for optimizing performance through advanced quantization techniques. The experiments were conducted on the following instruction-tuned LLMs and the model description and licenses are discussed in Table 6.
Model License Type Usage Restrictions
Llama 3 8B Research License Non-commercial, requires attribution, ethical guidelines
Phi 3 3.8B Open-source Broad use allowed, must include original license and notices
Mistral 7B Open-source Broad use allowed, must include original license and notices
Gemma 7B Open-source/Research Depending on license, may have non-commercial restrictions or broad use allowed
Table 6: License information for LLMs used in the experiments.
The selection of Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B offers a comprehensive evaluation set that spans different model sizes, architectures, and use cases, ensuring the HPF framework is robust and generalizable. Llama 3 8B excels in language understanding and generative tasks, achieving top-tier performance across benchmarks such as MMLU, HumanEval, GSM-8K, BoolQ, and CommonSenseQA. This model’s proficiency in diverse tasks, including complex mathematical reasoning and commonsense reasoning, underscores its high-quality text generation capabilities. Phi 3 3.8B, optimized for environments with limited computational resources, performs competitively on benchmarks like GSM-8K and BoolQ, making it ideal for real-time applications due to its efficiency and quick response times. Mistral 7B balances parameter count and performance, excelling in MMLU, CommonSenseQA, and HumanEval, showcasing versatility in language understanding and code-related tasks. Gemma 7B, specialized for particular domains, excels in niche areas like medical text understanding and technical documentation. While not always the top performer in general benchmarks, Gemma 7B’s specialization complements the general-purpose nature of the other LLMs, ensuring the HPF framework is applicable to a wide range of scenarios and tasks.
By including these LLMs, the effectiveness of the HPF is highlighted, demonstrating its ability to accommodate a wide range of scenarios and tasks, from general-purpose language understanding and generation to specialized domain-specific applications. This comprehensive evaluation ensures that HPF is a robust and generalizable framework in the landscape of LLM evaluation and prompting frameworks.
The quantization configuration for each of the instruction-tuned LLM involves reducing the precision of the LLM’s weights and computations, allowing for a more compact representation and faster processing. The LLMs were loaded in a 4-bit precision format which significantly reduces the memory footprint without compromising accuracy. By using half-precision floating point computations (float16), the LLMs maintain high computational efficiency. Additionally, the quantization type employed is NF4, which provides a balanced approach between model size and performance. The use of double quantization further refines this process, ensuring that the LLMs operate efficiently even on resource-constrained hardware.
The generation configuration is carefully designed to optimize the quality of text output during evaluation. This involves setting parameters that influence the behavior of the LLMs during the text generation process. For instance, the maximum number of tokens that can be generated in a single run is capped (1024 tokens) to ensure concise and relevant outputs. The temperature parameter was set to 0.6 to control the randomness of the predictions; lower temperature results in more deterministic and focused outputs. Top-p (p=0.9) sampling allows the LLMs to explore various possible continuations for a given input. Finally, the repetition penalty is applied to discourage the LLMs from repeating the same phrases or sentences, thus ensuring more coherent and varied text generation.
Appendix B Prompt Templates
B.1 Level 1: Role Prompting
Role prompting is the most basic level of interaction with a LLM, where the LLM is given a specific role or task to perform without additional context or examples. This technique relies solely on the initial instruction to guide the LLM’s response. For instance, if the LLM is asked to "act as a translator", it will attempt to translate text based on its training data. This method is straightforward but may lack depth and context, leading to less accurate or nuanced outputs. Table 7 presents the prompt templates utilized for Role prompting across all datasets in our experiments.
Dataset Prompt
BoolQ Based on the passage: ‘‘passage", answer True/False to the question: ‘‘question" as an Omniscient person.
CSQA Choose the answer: ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5" as a critical thinker.
IWSLT Translate ‘‘english text" to french as a Translator.
SamSum Summarize the Dialogue: ‘‘dialogue" as a Summarizer.
Table 7: Prompt templates of different datasets for Role Prompting.
B.2 Level 2: Zero-shot Chain-of-Thought Prompting
Zero-shot Chain-of-Thought prompting builds upon basic role prompting by asking the LLM to generate a reasoning process for a given task, even though it has not been explicitly trained on similar examples. In this approach, the LLM attempts to break down the problem and solve it step-by-step using its internal knowledge base. This technique leverages the LLM’s ability to infer and reason, improving the quality of the response by encouraging logical progression and coherence in the thought process. Table 8 presents the prompt templates utilized for Zero-CoT across all datasets in our experiments.
Dataset Prompt
BoolQ Based on the passage: ‘‘passage", answer True/False to the question: ‘‘question". Let’s think step by step.
CSQA Choose the answer: A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5". Let’s think step by step.
IWSLT Translate ‘‘english text" to french. Let’s think step by step.
SamSum Summarize the Dialogue: ‘‘dialogue". Let’s think step by step.
Table 8: Prompt templates of different datasets for Zero-shot Chain-of-Thought Prompting.
B.3 Level 3: Three-Shot Chain-of-Thought Prompting
Three-shot Chain-of-Thought prompting advances the zero-shot approach by providing the LLM with three examples of the task at hand, including the reasoning steps taken to arrive at the solution. These examples serve as a guide, helping the LLM understand the structure and logic required to solve similar problems. By seeing how previous tasks were approached and resolved, the LLM can better mimic the process, leading to more accurate and contextually appropriate responses. Table 9 presents the prompt templates utilized for 3-CoT across all datasets in our experiments.
Dataset Prompt
BoolQ Based on the passage: "passage1", answer True/False to the question: "question1". Answer: "answer1". Explanation: "explaination1". Based on the passage: "passage2", Answer True/False to the question: "question2". Answer: "answer2". Explanation: "explaination2". Based on the passage: "passage3", Answer True/False to the question: "question3". Answer: "answer3". Explanation: "explaination3". Based on the passage: "passage", Answer True/False to the question: "question".
CSQA Choose the answer: "question1",A. "option1-1",B. "option2-1",C. "option3-1",D. "option4-1",E. "option5-1", Answer: "answer1", Explanation: "explaination1". Choose the answer: "question2", A. "option1-2",B. "option2-2",C. "option3-2",D. "option4-2",E. "option5-2", Answer: "answer2", Explanation: "explainatio n2". Choose the answer: "question3", A. "option1-3",B. "option2-3",C. "option3-3",D. "option4-3",E. "option5-3", Answer: "answer3", Explanation: "explaination3". Choose the answer: "question", ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5".
IWSLT Translate "english text1" to French. French: "french text1". Translate "english text2" to French. French: "french text2". Translate "english text3" to French. French: "french text3". Translate "english text" to French.
SamSum Summarize the Dialogue: "dialogue1". Summary: "summary1". Summarize the Dialogue: "dialogue2". Summary: "summary2". Summarize the Dialogue: "dialogue3". Summary: "summary3". Summarize the Dialogue: "dialogue".
Table 9: Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting
B.4 Level 4: Least-to-Most Prompting
Least-to-most prompting is a sophisticated technique that involves gradually increasing the complexity of prompts presented to the LLM. The process begins with simpler, more manageable tasks and progressively moves to more complex challenges. This method helps the LLM build confidence and competence in solving problems, as it can leverage the insights gained from simpler prompts to tackle more difficult ones. It enhances the LLM’s ability to generalize from straightforward examples to more intricate scenarios. Table 10 presents the prompt templates utilized for Least-to-Most Prompting across all datasets in our experiments.
Dataset Prompt
BoolQ prompt 1: Summarize the main points of this passage: "passage". prompt 2: Analyze this question to identify its key components: "question".
prompt 3: Find the part of the passage that relates to this question: "question", Passage: "passage".
prompt 4: Based on the passage, what is the answer to this question: "question", Relevant Information: "previous response".
CSQA prompt 1: Analyze this question: "question".
prompt 2: Elaborate about each option for the question: "question", options: A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5".
prompt 3: Based on the analysis: "previous response", discard wrong answers among the options: A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5".
prompt 4: Choose the correct answer from the options: A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5".
IWSLT prompt 1: What is the main idea or theme of this text? "english text".
prompt 2: Identify and list the key phrases or terms in this text: "english text".
prompt 3: Translate the following key phrases into French: "previous response".
prompt 4: Translate "english text" into French, incorporating the translations of the key phrases: "previous response".
SamSum prompt 1: List the main points or key ideas present in this dialogue: "dialogue".
prompt 2: Elaborate on the following key points, providing additional details or context: "previous response".
prompt 3: Using the listed key points and their elaborations, draft a concise summary of this text: "dialogue".
prompt 4: Refine this draft summary to make it more concise and coherent, ensuring it captures the essence of the text: "dialogue".
Table 10: Prompt templates of different datasets for Least-to-Most Prompting
B.5 Level 5: Generated Knowledge Prompting
Generated Knowledge prompting represents one of the most complex prompting techniques present in the HPF, where the LLM not only addresses the given task but also incorporates additional relevant information to enhance its response. This method involves prompting another LLM to produce auxiliary knowledge that can be used to better understand and solve the problem. By creating a rich context and leveraging self-generated insights, the LLM can provide more detailed, accurate, and nuanced answers, showcasing the full extent of its capabilities. Table 11 presents the prompt templates utilized for Least-to-Most Prompting across all datasets in our experiments.
Dataset Prompt
BoolQ inference prompt: Based on the passage:"passage", answer True/False to the question: ’question’ using knowledge of the passage:"knowledge"
knowledge generation prompt: Generate Knowledge about the passage: "passage".
CSQA inference prompt: Choose the answer:"question", A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5" using knowledge of the question:"knowledge"
knowledge generation prompt: Generate Knowledge about the question: "question".
IWSLT inference prompt: Translate "english text": to French using definitions of the keywords:"knowledge"
knowledge generatiuon prompt: Generate definitions in french of each word in the text: "english text".
SamSum inference prompt: Summarize the Dialogue: "dialogue" using the interpretation of the dialogue:"knowledge"
knowledge generation prompt: Generate interpretation about the dialogue: "dialogue".
Table 11: Prompt templates of different datasets for Generated Knowledge Prompting
B.6 Adaptive HPF
The prompt-selector in adaptive HPF selects the prompting level based on the task complexity to address the task. Llama 3 8B serves as the prompt-selector in the experiments. The prompt template was meticulously designed to ensure maximum clarity, aiming to reduce hallucinations and select the most effective prompting strategy.
Prompt Template: Choose the most effective prompting strategy among five available strategies for the task. Begin with the lowest indexed strategy and progress to higher indexed strategies if the earlier ones are not effective. For a given task, the prompting strategies are:
  • Role Prompting: Defines a role for the model in solving the task.
  • Zero-shot Chain of Thought prompting: Stimulates reasoning and problem-solving by including the phrase ’Let’s think step by step’ without offering previous examples related to the task.
  • Three-shot Chain of Thought prompting: Offers three examples related to the task to guide the model’s reasoning process.
  • Least-to-most prompting: Uses a sequential method to derive essential insights from the task in order to solve it.
  • Generated Knowledge Prompting: Integration and application of external knowledge to accomplish the task. The external knowledge is generated using some other model based on the task.
Select only the index (do not provide the name) of the most effective prompting strategy.
Figure 8: HP-Scores across all datasets and LLMs for manual HPF and Adaptive HPF.
Appendix C Human Annotation Policy
To ensure the quality of HP-Scores scored by human experts for the datasets, we implemented a rigorous human annotation process. Human annotators were tasked with calculating the HP-Score for each sample in a given dataset. The HP-Score relative to human experts is a measure of the expert capability in addressing the task based on criteria of HPT. Each sample was scored on a scale from 1 (Lowest prompting level) to 5 (Highest prompting level) for the following criteria:
C.1 Criteria
  1. 1.
    Basic Understanding and Reproduction
    • This criterion evaluates the annotator’s ability to comprehend and accurately reproduce the content. A high score indicates clear and precise reproduction of the information is required for solving the task.
  2. 2.
    Understanding and Interpretation
    • This criterion assesses the depth of understanding and the ability to interpret the information correctly. A high score indicates a detailed and precise understanding of the task is necessary to solve the task.
  3. 3.
    Analysis and Reasoning
    • This criterion measures the ability to analyze the information and apply logical reasoning. A high score demonstrates the need to analyze the information to solve the task.
  4. 4.
    Application of Knowledge and Execution
    • This criterion evaluates the practical application of knowledge and the execution of tasks based on the relevant knowledge. A high score indicates effective application and the requirement of knowledge to solve the task.
C.2 Scoring Process
Each sample of every dataset was scored individually by the annotators on a scale from 1 to 5, where:
  • 1 = Poor
  • 2 = Fair
  • 3 = Good
  • 4 = Very Good
  • 5 = Excellent
The final HP-Score for each dataset was calculated by taking the mean of the values from these four criteria. This approach ensures a comprehensive evaluation of each sample by considering multiple performance criteria.
Appendix D Additional Results
D.1 HP-Scores Interpretation
Figure 8 visualizes HP-Scores for different datasets across various LLMs, comparing manual and adaptive frameworks. Here’s a detailed summary:
For the BoolQ dataset, the manual HPF HP-Scores are generally lower across all LLMs compared to the Adaptive HPF. The adaptive scores are significantly higher, indicating better performance with the manual HPF for solving BoolQ.
For the CSQA dataset, manual HPF HP-Scores are moderate, slightly higher than those for BoolQ. However, similar to BoolQ, the Adaptive HPF HP-Scores are consistently higher, again showing the superiority of the manual HPF over adaptive HPF in solving CSQA.
In comparison to the other datasets, the manual HPF HP-Scores for the IWSLT (0.15) and IWSLT (0.20) datasets both show a significant increase. This increase is particularly noticeable for LLMs such as Phi 3 3.8B and Gemma 7B, which indicate a lack of ability to handle IWSLT. With comparable evaluation scores to manual HPF, which suggests no improvement in adaptive HPF, the adaptive HPF HP-Scores continue to be consistently high.
For both SamSum (0.15) and SamSum (0.20) datasets, the manual HPF HP-Scores are notably high for LLMs such as Phi 3 3.8B and Gemma 7B. Adaptive HPF HP-Scores are consistently high compared to manual HPF HP-Scores.
As evidenced by their lower HP-Scores, manual HPF routinely outperforms adaptive HPF in all tasks and LLMs, including BoolQ, CSQA, and SamSum, where it can outperform human experts. With the exception of solving SamSum (0.15) by Mistral 7B in Adaptive HPF, no LLM was able to complete any task with a HP-Score of less than 5, suggesting the need for additional prompting strategies in HPF or improved framework to avoid hallucinations in selecting prompting level by prompt-selector.
D.2 Hallucination in Adaptive HPF
Hallucinations in prompt-selector refer to instances where the LLM generates incorrect or misleading prompting levels, or nonsensical information that is not supported by HPF. These hallucinations can occur across various tasks, including question answering, multiple-choice questions, translation, and summarization. Below, we discuss the occurrence of hallucinations in the Llama 3 8B which was utilized as prompt-selector in our experiments across different tasks and iterations, highlighting how the LLM’s performance varies and where it tends to hallucinate.
As shown in Figure 9, for the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations (represented by ’…’). However, by the fourth iteration, Llama 3 8B manages to answer correctly at Level 2. For the CSQA task, prompt-selector exhibits hallucinations initially, shown by Level 4 and Level 0 (not included in HPF) responses. Eventually, it corrects itself by the third iteration, providing the correct answer at Level 2. For the IWSLT task, prompt-selector demonstrates a consistent pattern of hallucinations across multiple iterations. Even though Llama 3 8B attempts the translation at Level 2 multiple times, it ultimately fails to provide a correct translation, indicating a persistent hallucination. For the SamSum task, prompt-selector shows initial hallucinations in the first three iterations (Level 4). However, by the fourth and fifth iterations, the prompt-selector starts producing lower levels, finally, Llama 3 8B achieving correct answer at Level 2 in the last iteration .
The results indicate that Llama 3 8B exhibits hallucinations in selecting prompting levels across various tasks and iterations resulting in higher HP-Scores for Adaptive HPF, with performance varying significantly. While the LLM can eventually produce correct answers, as seen in the BoolQ and SamSum tasks, it often requires multiple attempts and may still fail in tasks like IWSLT translation.
Figure 9: Hallucination of prompt-selector in choosing the most effective prompting level to solve corresponding tasks by Llama 3 8B.
Appendix E Additional Experiments
E.1 Increasing Threshold for IWSLT and SamSum Datasets
When comparing the performance of various LLMs, Llama 3-8B stands out for its superior results on the IWSLT and SamSum tasks, consistently performing better in both manual and adaptive frameworks. At a threshold of 0.2, the performance of all LLMs reached saturation for these tasks, indicating that increasing the threshold further did not yield improvements. Both adaptive and manual frameworks struggled with thresholds higher than 0.2, failing to meet or exceed this benchmark for most samples within their iterations. Even the most capable LLM, Llama 3-8B, was unable to address the majority of samples across any prompting level within the HPF. This demonstrates the limitations of current HPFs at higher thresholds and highlights the ongoing challenges in achieving better performance in LLMs.
E.2 Adaptive HPF with Varying Prompt-selectors
Experiments demonstrated that Adaptive HPF is most effective when Llama 3 8B is used as a prompt-selector compared to LLMs, attributed to its ability to produce the lowest number of hallucinations when selecting the prompting level within the iterations of the adaptive HPF. Conversely, Gemma 7B, which exhibited numerous hallucinations across most tasks, was found to be the least effective adaptive HPF. Consequently, the choice of the prompt-selector emerges as a crucial factor influencing the efficiency of the adaptive HPF.

### v4

@@
-Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models
+Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models Aligned with Human Cognitive Principles
 Abstract
-Assessing the effectiveness of large language models (LLMs) in addressing diverse tasks is essential for comprehending their strengths and weaknesses.
-Conventional evaluation techniques typically apply a single prompting strategy uniformly across datasets, not considering the varying degrees of task complexity.
-We introduce the Hierarchical Prompting Taxonomy (HPT), a taxonomy that employs a Hierarchical Prompt Framework (HPF) composed of five unique prompting strategies, arranged from the simplest to the most complex, to assess LLMs more precisely and to offer a clearer perspective.
-This taxonomy assigns a score, called the Hierarchical Prompting Score (HP-Score), to datasets as well as LLMs based on the rules of the taxonomy, providing a nuanced understanding of their ability to solve diverse tasks and offering a universal measure of task complexity.
-Additionally, we introduce the Adaptive Hierarchical Prompt framework, which automates the selection of appropriate prompting strategies for each task.
-This study compares manual and adaptive hierarchical prompt frameworks using four instruction-tuned LLMs, namely Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B, across four datasets: BoolQ, CommonSenseQA (CSQA), IWSLT-2017 en-fr (IWSLT), and SamSum.
-Experiments demonstrate the effectiveness of HPT, providing a reliable way to compare different tasks and LLM capabilities.
-This paper leads to the development of a universal evaluation metric that can be used to evaluate both the complexity of the datasets and the capabilities of LLMs.
-The implementation of both manual HPF and adaptive HPF is publicly available [^1] GitHub repository consisting of the code implementation of this work: https://github.com/devichand579/HPT.
-Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models
-Devichand Budagam1, Sankalp KJ2, Ashutosh Kumar3, Vinija Jain4, 5thanks: Work does not relate to position at Amazon, Aman Chadha4, 5thanks: Work does not relate to position at Amazon 1Indian Institute of Technology Kharagpur, India, 2AI Institute, University of South Carolina, USA 3Rochester Institute of Technology, USA, 4Amazon GenAI, USA, 5Stanford University, USA
+Assessing the effectiveness of large language models (LLMs) in performing different tasks is crucial for understanding their strengths and weaknesses.
+This paper presents Hierarchical Prompting Taxonomy (HPT), grounded on human cognitive principles and designed to assess LLMs by examining the cognitive demands of various tasks.
+The HPT utilizes the Hierarchical Prompting Framework (HPF), which structures five unique prompting strategies in a hierarchical order based on their cognitive requirement on LLMs when compared to human mental capabilities.
+It assesses the complexity of tasks with the Hierarchical Prompting Index (HPI), which demonstrates the cognitive competencies of LLMs across diverse datasets and offers insights into the cognitive demands that datasets place on different LLMs.
+This approach enables a comprehensive evaluation of an LLM’s problem-solving abilities and the intricacy of a dataset, offering a standardized metric for task complexity.
+Extensive experiments with multiple datasets and LLMs show that HPF enhances LLM performance by 2%-63% compared to baseline performance, with GSM8k being the most cognitively complex task among reasoning and coding tasks with an average HPI of 3.20 confirming the effectiveness of HPT.
+To support future research and reproducibility in this domain, the implementations of HPT and HPF are available here.
+Codehttps://github.com/devichand579/HPT
 1 Introduction
-Large language models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements across a wide range of applications.
-However, evaluating the capabilities of LLMs across various datasets remains a challenging task.
-Traditional prompting strategies, which apply the same prompt to all dataset samples regardless of task complexity, often result in sub-optimal evaluations.
+Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements in a wide range of applications.
+Conventional evaluation frameworks often apply a standard prompting approach to assess different LLMs, regardless of the complexity of the task, which may result in biased and suboptimal outcomes.
+Moreover, applying the same prompting approach across all samples within a dataset without considering each sample’s relative complexity adds to the unfair situation.
+To achieve a more balanced evaluation framework, it is essential to account for both the task-solving ability of LLMs and the varying cognitive complexities of the dataset samples.
 This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of task complexity.
-Figure 1: Hierarchical Prompt Framework (HPF) consists of five distinct prompting strategies, each tailored to different levels of task complexity, ensuring the appropriate prompt is selected for the given task.
+Figure 1: Hierarchical Prompting Framework includes five distinct prompting strategies, each designed for different levels of task complexity to ensure the appropriate prompt is selected for the given task.
+A $\checkmark$ indicates task completion, while a $\times$ signifies task incompletion.
+Within this study, complexity is defined as the cognitive demands associated with solving a task or the cognitive load introduced by a prompting strategy on LLMs.
+Henceforth, the term ”complexity” will be applied solely in this context.
+Task complexity, within the realm of human cognition, pertains to the cognitive requirements that a task imposes, which includes the diverse levels of mental effort necessary for processing, analyzing, and synthesizing information.
+According to Sweller (1988), tasks become more complex as they require greater cognitive resources, engaging working memory in more demanding processes such as reasoning and problem-solving.
+Similarly, Anderson et al.
+(2014) highlights that human cognitive abilities span a continuum from basic recall to higher-order thinking, with increasing difficulty correlating to tasks that demand analysis, synthesis, and evaluation.
+When applied to LLMs, the complexity of prompting strategies can be systematically evaluated by mapping them onto this human cognitive hierarchy.
+This alignment allows for an assessment of how LLMs perform tasks that reflect varying degrees of cognitive load, thereby providing a structured framework for understanding the cognitive demands associated with various tasks.
+By imposing this cognitive complexity framework on LLMs, this paper establishes a universal evaluation method, grounded in human cognitive principles, that enables more precise comparisons of model performance across tasks with varying levels of difficulty.
+This paper introduces the HPT, a set of rules that maps the human cognitive principles for assessing the complexity of different prompting strategies.
+It employs the HPF shown in Figure 1, a prompt selection framework that selects the prompt with the optimal cognitive load on LLM required in solving the task.
+HPF enhances the interaction with LLMs, and improves performance across various tasks by ensuring prompts resonate with human cognitive principles.
+The main contributions of this paper are as follows:
+  1. 1.
+    Hierarchical Prompting Taxonomy (HPT): The paper proposes HPT, a set of rules that maps prompting strategies onto human cognitive principles, thereby enabling a universal measure of task complexity for LLMs.
+  2. 2.
+    Hierarchical Prompting Framework (HPF): The HPF is a framework designed to select the most effective prompt from five distinct prompting strategies, which optimizes the cognitive load on LLMs during task completion.
+This framework not only facilitates a more accurate evaluation of LLMs but also enhances their performance, delivering more transparent insights.
+  3. 3.
+    Hierarchical Prompting Index (HPI): HPI [^1] HPI can be quantitatively assessed to analyze the cognitive abilities of an LLM and the cognitive demands imposed by datasets on LLMs, as both factors are interchangeably related to the complexity of tasks.
+quantitatively evaluates the task complexity of LLMs across various datasets, offering insights into the cognitive demands that each task imposes on different LLMs.
+HPF can be effectively compared to an ”Open Book” examination as shown in Figure 2, where questions represent tasks and textbooks serve as prompting strategies.
+In this analogy, the exam questions vary in complexity, from simple factual recall to intricate analytical problems, analogous to the tasks in HPT that are evaluated based on their cognitive demands.
+Similarly, textbooks provide structured guidance for solving these questions, much like the HPF organizes prompts in increasing levels of complexity to support LLMs.
+For example, a straightforward glossary lookup corresponds to a low-complexity task, while solving a multi-step analytical problem requiring synthesis of concepts represents a high-complexity task.
+The effort a student invests in answering a question mirrors the HPI, which measures the cognitive load placed on the LLM.
+Just as students perform better with structured resources like textbooks, LLMs improve with well-designed hierarchical prompting strategies, enabling them to tackle progressively complex tasks effectively.
+Figure 2: Analogical framework comparing the HPF with ”Open Book” examination methodology.
+The diagram illustrates how HPF components (below) mirror traditional educational assessment elements (above), with parallel relationships between task complexity levels, resource utilization (prompts/textbooks), and performance metrics (HPI/student effort).
+This comparison demonstrates how LLM task complexity scales similarly to educational assessment complexity, from simple lookup tasks to complex synthesis problems
+The remainder of the paper is structured as follows: Section 2 reviews the related work on prompting and evaluation in LLMs.
+Section 3 details the HPT and its associated frameworks.
+Section 4 outlines the experimental setup, results, and ablation studies.
+Section 5 concludes the paper.
+Section 6 discusses the ethical impact of the work.
+2 Related Work
+The advent of LLMs has revolutionized NLP by demonstrating significant improvements in few-shot and zero-shot learning capabilities.
+Brown et al.
+(2020) introduced GPT-3, a 175 billion parameter autoregressive model, showcasing its ability to perform a wide range of tasks such as question-answering, reading comprehension, translation, and natural language inference without fine-tuning.
+This study highlighted the potential of very large models for in-context learning while also identifying limitations in commonsense reasoning and specific comprehension tasks.
+Similarly, Liu et al.
+(2021) surveyed prompt-based learning, emphasizing the role of prompt engineering in leveraging pre-trained models for few-shot and zero-shot adaptation to new tasks with minimal labeled data.
+2.1 Prompt Engineering
 Prompting plays a vital role in unlocking the full potential of LLMs.
@@
 Recent research has explored various approaches to prompting and reasoning evaluation in LLMs.
-Chain-of-Thought (CoT) prompting Wei et al.
-(2022b) elicits step-by-step reasoning, improving performance on complex tasks.
-Specializing smaller models Fu et al.
-(2023) and using large models as reasoning teachers Ho et al.
-(2022) have demonstrated the potential for enhancing reasoning capabilities.
-Progressive-hint prompting Zheng et al.
-(2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses.
-Metacognitive prompting Wang and Zhao (2024) incorporates self-aware evaluations to enhance understanding abilities.
-We introduce the Hierarchical Prompting Taxonomy (HPT), a set of rules that employs a Hierarchical Prompt Framework (HPF) shown in Figure 1, a structured evaluation method addressing the need for more sophisticated assessment of LLM task-solving capabilities.
-This framework employs five distinct prompting strategies, each tailored to different levels of task complexity.
-By assigning a score called the HP-Score based on the LLM’s performance across these levels, this work provides a comprehensive assessment of its ability to solve diverse tasks.
-Furthermore, the taxonomy incorporates an Adaptive Hierarchical Prompt framework, enabling an LLM, called prompt-selector to dynamically select the appropriate prompting strategy based on a given task complexity.
-This adaptive approach automates the evaluation process, also provides a better understanding of the judgment skills of LLMs in determining task complexity.
-The main contributions of this paper are as follows:
-  1. 1.
-    We propose the Hierarchical Prompting Taxonomy (HPT), a set of rules designed to establish a universal measure of task complexity for both datasets and LLMs.
-  2. 2.
-    Based on the rules of the HPT, this work introduces a set of five unique prompting strategies, arranged from simplest to most complex, to provide a more precise assessment of LLMs and offer clearer insights.
-  3. 3.
-    We introduce the Adaptive HPF that allows us to examine the judgement skills of LLMs by automating the evaluation process and dynamically selecting the right prompting approach based on task complexity.
-The rest of the paper is organized as follows: Section 2 provides an overview of related work on prompting and evaluation in LLMs.
-Section 3 presents HPT and the frameworks associated with it in detail.
-Section 4 describes the experimental setup and results.
-Section 5 concludes the paper, Section 6 discusses the limitations and finally, Section 7 discusses the ethical impact of this work.
-2 Related Work
-The advent of LLMs has revolutionized NLP by demonstrating significant improvements in few-shot and zero-shot learning capabilities.
-Brown et al.
-(2020) introduced GPT-3, a 175 billion parameter autoregressive model, showcasing its ability to perform a wide range of tasks such as question-answering, reading comprehension, translation, and natural language inference without fine-tuning.
-This study highlighted the potential of very large models for in-context learning, while also identifying limitations in commonsense reasoning and specific comprehension tasks.
-Similarly, Liu et al.
-(2021) surveyed prompt-based learning, emphasizing the role of prompt engineering in leveraging pre-trained models for few-shot and zero-shot adaptation to new tasks with minimal labeled data.
+Chain-of-Thought (CoT) prompting (Wei et al.
+2022b) elicits step-by-step reasoning, improving performance on complex tasks.
+Specializing smaller models (Fu et al.
+2023) and using large models as reasoning teachers (Ho, Schmid, and Yun 2022) have demonstrated the potential for enhancing reasoning capabilities.
 Emergent abilities in LLMs, which appear suddenly at certain scale thresholds, have also been a topic of interest.
@@
 (2022) demonstrated that LLMs could exhibit multi-step reasoning capabilities in a zero-shot setting by simply modifying the prompt structure, thus highlighting their potential as general reasoning engines.
-Further exploring reasoning capabilities, Qiao et al.
-(2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning.
-Several studies have focused on improving the interaction and decision-making abilities of LLMs through innovative prompting techniques.
-Lo et al.
-(2023) proposed ASH (Actor-Summarizer-Hierarchical) prompting, which enhances LLM performance on complex web navigation tasks by decomposing the process into summarization and action prediction.
-This hierarchical approach significantly outperformed previous methods on tasks with lengthy trajectories.
+Yao et al.
+(2023) introduced the Tree-of-Thoughts framework, enabling LLMs to deliberate over coherent text units and perform heuristic searches for complex reasoning tasks.
+This approach generalizes over chain-of-thought prompting and has shown significant performance improvements in tasks requiring planning and search, such as creative writing and problem-solving games.
 Kong et al.
 (2024) introduced role-play prompting to improve zero-shot reasoning by constructing role-immersion interactions, which implicitly trigger chain-of-thought processes and enhance performance across diverse reasoning benchmarks.
-The robustness and safety of LLM outputs have also been addressed in recent research.
+Progressive-hint prompting (Zheng et al.
+2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses.
+Metacognitive prompting (Wang and Zhao 2024) incorporates self-aware evaluations to enhance understanding abilities.
+These works collectively highlight the advancements in leveraging LLMs through innovative prompting techniques, addressing their emergent abilities, reasoning capabilities, interaction strategies, robustness, and evaluation methodologies.
+Despite significant advancements, the current LLM research reveals several limitations, particularly in terms of prompt design, handling complex reasoning tasks, and evaluating model performance across diverse scenarios.
+While promising, the emergent abilities of LLMs often lack predictability and control, and the robustness of these LLMs in the face of misleading prompts remains a concern.
+2.2 Prompt Optimization and Selection
+The challenge of optimizing prompts for LLMs has been addressed in several key studies, each contributing unique methodologies to enhance model performance and efficiency.
+Shen et al.
+(2023) introduce PFLAT, a metric utilizing flatness regularization to quantify prompt utility, which leads to improved results in classification tasks.
+Do et al.
+(2024) propose a structured three-step methodology that contains data clustering, prompt generation, and evaluation, effectively balancing generality and specificity in prompt selection.
+ProTeGi (Pryzant et al.
+2023) offers a non-parametric approach inspired by gradient descent, leveraging natural language ”gradients” to iteratively refine prompts.
 Wang et al.
-(2023) developed INDUST, a benchmark to evaluate LLM resistance to inductive instructions containing false premises, showing that self-critique prompting can help models handle such challenging instructions more effectively.
-Additionally, Chen et al.
-(2024) provided a comprehensive survey on prompt engineering techniques aimed at reducing hallucinations and enhancing the reliability of LLM outputs through advanced methods like chain-of-thought and tree-of-thought prompting.
+(2024) present PromISe, which transforms prompt optimization into an explicit chain of thought, employing self-introspection and refinement techniques.
+Zhou et al.
+(2023b) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
+These studies aim to automate prompt design, moving away from traditional manual trial-and-error methods while emphasizing efficiency and scalability across various tasks and models.
+They report significant improvements in LLMs performance, with enhancements ranging from 5% to 31% across different benchmarks.
+This body of work underscores the increasing importance of prompt optimization and selection in unlocking the potential of LLMs and points toward future research avenues, such as exploring theoretical foundations, integrating multiple optimization techniques, and distinguishing between task-specific and general-purpose strategies.
+2.3 Evaluation Benchmarks
 To facilitate the evaluation and understanding of LLM capabilities, Zhu et al.
@@
 This modular and extensible tool aims to support collaborative research and advance the comprehension of LLM strengths and weaknesses.
-Zhou et al.
-(2023b) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
-Chang et al.
-(2022) proposed SpeechPrompt, a framework that reformulates tasks as conditional generation problems for pre-trained generative spoken language models.
-This method demonstrated competitive performance in speech classification with fewer trainable parameters and potential in more challenging tasks like automatic speech recognition.
-Lastly, Yao et al.
-(2023) introduced the Tree-of-Thoughts framework, enabling LLMs to deliberate over coherent text units and perform heuristic searches for complex reasoning tasks.
-This approach generalizes over chain-of-thought prompting and has shown significant performance improvements in tasks requiring planning and search, such as creative writing and problem-solving games.
-These works collectively highlight the advancements in leveraging LLMs through innovative prompting techniques, addressing their emergent abilities, reasoning capabilities, interaction strategies, robustness, and evaluation methodologies.
-Despite significant advancements, the current research on LLMs reveals several limitations, particularly in terms of prompt design, handling complex reasoning tasks, and evaluating model performance across diverse scenarios.
-The emergent abilities of LLMs, while promising, often lack predictability and control, and the robustness of these LLMs in the face of misleading prompts remains a concern.
-To address these challenges, we propose the HPT and HPF for evaluating different LLMs.
-This approach aims to systematically categorize and assess the effectiveness of various prompting techniques based on task complexity, enabling a more nuanced understanding of LLM behaviors and performance.
-By introducing a structured framework for evaluation, this work seeks to enhance the reliability and applicability of LLMs, facilitating their deployment in increasingly complex and dynamic environments.
-3 Methodology
-Figure 2: Hierarchical Prompting Taxonomy (HPT): A taxonomy designed to assess task complexity based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
-3.1 Datasets and Models
-Experiments were performed on BoolQ Clark et al.
-(2019), CommonSenseQA (CSQA) Talmor et al.
-(2019), IWSLT-2017 en-fr (IWSLT) Cettolo et al.
-(2017), and SamSum Gliwa et al.
-(2019).
-BoolQ consists of approximately 16,000 True/False questions with corresponding passages from Wikipedia, designed for binary classification in reading comprehension.
-CSQA contains around 12,000 multiple-choice questions sourced from ConceptNet, aimed at evaluating model’s commonsense reasoning abilities.
-IWSLT is a parallel corpus with tens of thousands of English-French sentence pairs from TED Talks, used for machine translation tasks.
-SamSum features around 16,000 human-generated chat logs with summaries, facilitating dialogue summarization model development.
-On the aforementioned datasets, the performance of the instruction-tuned LLMs Llama 3 8B AI@Meta (2024), Phi 3 3.8B Abdin et al.
-(2024), Mistral 7B Jiang et al.
-(2023), and Gemma 7B Team et al.
-(2024) were evaluated on their validation splits, with the exception of SamSum, which utilized the test split and nearly 2300 hours of computation was used for all experiments on the V100 GPU.
-3.2 Hierarchical Prompting Taxonomy (HPT)
-The proposed HPT shown in Figure 2 is founded on principles that are crucial for humans to address the majority of reasoning and interpretation tasks, taking into account the following criteria:
-  • Basic Recall and Reproduction: Remember and reproduce factual information without interpretation or analysis.
-  • Understanding and Interpretation: Comprehend and explain the meaning of information, summarizing or clarifying content.
-  • Analysis and Reasoning: Break down complex information, understand relationships, and solve problems using logical reasoning.
-  • Application of Knowledge and Execution: Apply knowledge in practical situations, execute multi-step processes, and solve complex tasks.
-3.3 Hierarchical Prompt Framework (HPF)
-The HPT employs an HPF that comprises five different prompting strategies.
-The complexity levels of these prompting strategies within the framework are determined by the taxonomy criteria, which judges actions such as the requirement for external knowledge, the capacity to solve subproblems, and the capacity to provide step-by-step reasoning that takes the task’s complexity into account.
-The five different prompting strategies that make up the HPF are :
-  • Role Prompting Kong et al. (2024): Prompts that define a role for the LLM in solving the task.
-  • Zero-Shot Chain-of-Thought Prompting (Zero-CoT) Kojima et al.
-(2022): Prompts that use the phrase "Let’s think step by step" without providing prior examples to encourage critical thinking and problem-solving.
-  • Three-Shot Chain-of-Thought Prompting (3-CoT) Wei et al. (2022b): Prompts that offer three examples to guide the LLM’s reasoning process.
-  • Least-to-Most Prompting Zhou et al. (2023a): Prompts that sequentially breakdown the task into sub problems to derive essential insights from the task in order to solve it.
-  • Generated Knowledge Prompting (GKP) Liu et al.
-(2022): Prompts that require the integration of external knowledge to accomplish the task.
-Llama 3 8B was used to generate the necessary knowledge.
-Unlike the original study, the knowledge prediction with the highest confidence score was selected among three possible predictions.
-The Manual HPF systematically iterates through the predefined levels of prompting strategies to solve a given task using the language models.
-This process involves starting with the simplest prompts and gradually progressing to more complex ones based on the prompt hierarchy levels until the task is successfully addressed.
-By manually traversing each level, it ensures that every potential strategy is explored and allowing the LLM to identify the most effective approach for the task.
-A prompt-selector determines which prompting strategy is most suited for a given task in the adaptive HPF.
-This allows us to examine how LLMs may be utilised to assess their abilities to do different tasks while also automating the evaluation process.
-3.4 Hierarchical Prompting Score (HP-Score)
-The HP-Score is an evaluation metric for assessing the complexity of a task in comparison to various categories of problem-solving agents, such as humans and LLMs, which is heavily influenced by HPT rules.
-A task’s lower HP-Score indicates that the corresponding agent is more capable of solving the task.
-3.4.1 HP-Score for Datasets
-Datasets are evaluated by scoring each sample of a representative set of 5% of the original evaluation dataset on a scale from 1 (lowest) to the number of prompting levels (highest) existing in HPF for each of the four HPT rules.
-HP-Score for datasets is computed as:
-$\text{hp}_{\text{j}}=\frac{1}{4}\sum_{i=1}^{4}a_{i}$ (1)
+Further exploring reasoning capabilities, Qiao et al.
+(2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning.
+(Wang et al.
+2021) introduced a multi-task Benchmark for robustness Evaluation of LLMs extends the original GLUE (Wang et al.
+2018) benchmark to assess model robustness against adversarial inputs.
+It incorporates perturbed versions of existing GLUE tasks, such as paraphrasing, negation, and noise, to test models’ abilities with challenging data.
+The study highlights that despite their success on clean datasets, state-of-the-art models often struggle with adversarial examples, underscoring the importance of robustness evaluations in model development.
+3 Hierarchical Prompting Taxonomy
+Figure 3: Hierarchical Prompting Taxonomy: A taxonomy designed to assess the complexity of prompting strategies based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
+3.1 Governing Rules
+Figure 3 illustrates the HPT, a taxonomy that systematically reflects human cognitive functions as outlined in Bloom (1956).
+Each rule embodies complex cognitive processes based on established principles from learning and psychology.
+  1. 1.
+    Basic Recall and Reproduction: This reflects the fundamental cognitive process of remembering and reproducing factual information without analysis or interpretation, which involves mere recognition or retrieval of knowledge from memory (Anderson et al.
+2014).
+  2. 2.
+    Understanding and Interpretation: This corresponds to the second cognitive rule of (Bloom 1956), where individuals must not only recall information but also explain it in their own words, summarize key points or clarify the meaning of content.
+This rule demands an intermediate cognitive load involving information processing rather than retrieving it.
+  3. 3.
+    Analysis and Reasoning: This aligns with the analysis stage of (Bloom 1956), which involves higher cognitive functions such as comparison, contrast, and deep understanding of the underlying principles.
+It is more complex than mere understanding because it requires examining structure and identifying patterns and connections.
+  4. 4.
+    Application of Knowledge and Execution: This mirrors the application and evaluation stages of (Bloom 1956), where individuals must not only understand and analyze but also use knowledge to perform multi-step tasks, solve complex problems, and execute decisions.
+It represents the most cognitively complex tasks, which require synthesis of information and practical decision-making, highlighting the critical leap from understanding theory to executing it in practice.
+In HPT, the progression from basic recall to application of knowledge reflects increasing cognitive complexity, consistent with educational and cognitive frameworks, where more advanced cognitive processes build on foundational ones, demanding deeper engagement and mental effort.
+3.2 Hierarchical Prompting Framework
+The HPF consists of five prompting strategies, each assigned a complexity level.
+These levels are determined by the degree to which the strategies are shaped by the four principles of the HPT.
+The complexity levels of the prompting strategies are assigned based on human assessment of their relative cognitive loads over a set of 7 different tasks, guaranteeing that the cognitive abilities of LLMs are in harmony with those of humans.
+This approach enables the assessment of tasks in terms of their complexity and the cognitive load they impose on both humans and LLMs by utilizing HPI.
+Section 4.4 examines the hierarchical structure of the HPF in conjunction with the LLM-as-a-Judge framework, validating that the cognitive demands on LLMs can be aligned with those of humans.
+The set of five prompting strategies were chosen from a diverse range of existing strategies to populate the framework, guided by a human judgment policy, prioritizing comprehensiveness in cognitive demands rather than the sheer number of strategies.
+See Appendix A for more details.
+Consequently, the HPF can be replicated or expanded with other relevant prompting strategies that exhibit similar cognitive demands, making the framework adaptable.
+The five prompting strategies, listed from least to most complex, are as follows:
+  1. 1.
+    Role Prompting (Kong et al.
+2024): Prompts that specify the role of the LLM in task resolution represent the lowest level of complexity within the framework, as they exhibit minimal influence from the four principles of the HPT.
+  2. 2.
+    Zero-Shot Chain-of-Thought Prompting (Zero-CoT) (Kojima et al.
+2022): Prompts that use the phrase “Let’s think step by step” without providing prior examples aim to encourage critical thinking and problem-solving.
+This strategy implicitly activates chain-of-thought reasoning, demonstrating moderate influence from rule 3 of the HPT, while exerting a weaker influence from the other rules.
+  3. 3.
+    Three-Shot Chain-of-Thought Prompting (3-CoT) (Wei et al.
+2022b): Prompts that present three examples to guide the LLM’s reasoning process exert a strong influence from rules 1 and 2, requiring recall and interpretation of the examples to complete the task.
+This strategy also demonstrates moderate influence from rule 3, increasing its cognitive demands.
+  4. 4.
+    Least-to-Most Prompting (Zhou et al.
+2023a): A sequence of prompts that breaks the task into sub-problems exerts significant influence from all three rules 1, 2, and 3.
+This approach requires recalling prior prompts, interpreting previous responses, and analyzing them to effectively solve the task, resulting in a highly cognitively demanding strategy.
+  5. 5.
+    Generated Knowledge Prompting (GKP) (Liu et al.
+2022): Prompts that require integrating external knowledge to generate relevant information represent the most complex and cognitively demanding strategy.
+This approach is strongly influenced by rules 2, 3, and 4, as it involves correlating knowledge with the prompt and applying and analyzing external information, making it the most cognitively demanding within the HPT framework.
+In the experiments, Llama-3 8B is used to generate external knowledge.
+3.3 Hierarchical Prompting Index
+HPI is an evaluation metric for assessing the task complexity of LLMs over different datasets, which is influenced by the HPT rules.
+A lower HPI for a dataset suggests that the corresponding LLM is more adept at solving the task with fewer cognition processes.
+For each dataset instance, we begin with the least complex prompting strategy and progressively move through the HPF prompting strategies until the instance is resolved.
+The HPI corresponds to the complexity level of the prompting strategy where the LLM first tackles the instance.
+Algorithm 1 HPI Metric
 […]
-where $a_{i}$ represents the individual score for each rule of HPT, $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$ and $n$ denotes the total number of samples in the representative set.
-The HP-Scores assigned by experts in the domains related to the datasets used in the experiments are displayed in Table 1.
-CSQA is the most complex task relative to humans, with the highest HP-Score among the four datasets.
-Dataset Evaluation Set Size Representative Set Size HP-Score
+for sample $i$ in the evaluation dataset do
+     for level $x$ in the HPF do
+         if LLM resolves the task then
+[…]
+              break
+         end if
+     end for
+     if LLM failed to resolve the task then
+[…]
+     end if
+end for
+$\texttt{HPI}=\frac{1}{n}\sum_{j=1}^{n}\texttt{HPI\_List}[j]$
+$m$ is the total number of levels in the HPF, and $n$ is the total number of samples in the evaluation dataset.
+$\texttt{HPI}_{Dataset}$ represents the penalty introduced into the framework by human assessments.
+For further details on human annotation, see Appendix A.
+4 Results
+4.1 Experimental Setup
+Datasets
+The experiments utilized a diverse set of datasets, including MMLU, GSM8k, HumanEval, BoolQ, CSQA, SamSum, and IWSLT en-fr covering areas such as reasoning, coding, mathematics, question-answering, summarization, and machine translation, to evaluate the framework’s robustness and applicability.
+For further details on evaluation dataset sizes, see Appendix A.
+Reasoning: MMLU (Hendrycks et al.
+2021) includes multiple-choice questions across 57 subjects, covering areas like humanities, social sciences, physical sciences, basic mathematics, U.S.
+history, computer science, and law.
+CommonSenseQA (CSQA) (Talmor et al.
+2019) contains  12,000 questions to assess commonsense reasoning.
+Coding: HumanEval (Chen et al. 2021a) features 164 coding challenges, each with a function signature, docstring, body, and unit tests, designed to avoid training data overlap with LLMs.
+Mathematics: Grade School Math 8K (GSM8k) (Cobbe et al. 2021) comprises 8.5K diverse math problems for multi-step reasoning, focusing on basic arithmetic and early Algebra.
+Question-Answering: BoolQ (Clark et al. 2019) consists of  16,000 True/False questions based on Wikipedia passages for binary reading comprehension.
+Summarization: SamSum (Gliwa et al. 2019) features  16,000 human-generated chat logs with summaries for dialogue summarization.
+Machine Translation: IWSLT-2017 en-fr (IWSLT) (Cettolo et al. 2017) is a parallel corpus with thousands of English-French sentence pairs from TED Talks for translation tasks.
+Large Language Models
+For the evaluation, LLMs with parameter sizes ranging from 7 billion to 12 billion from top open-source models and top proprietary models were selected to determine the effectiveness of the proposed framework across varied parameter scales and architectures.
+Proprietary Models: GPT-4o (OpenAI 2024), Claude 3.5 Sonnet (Anthropic 2024)
+Open-Source models: Gemma 7B (Team et al. 2024a), Mistral 7B (Jiang et al. 2023), Llama-3 8B (AI@Meta 2024), Gemma-2 9B (Team et al. 2024b) and Mistral-Nemo 12B (Mistral AI and NVIDIA 2024).
+Additional Evaluation Metrics
+  • Coding: The Pass@k (Chen et al. 2021b) metric measures the probability of at least one correct solution among the top k outputs, used for evaluating code generation.
+  • Summarization: ROUGE-L (Lin 2004) evaluates the longest common subsequence between generated text and reference, focusing on sequence-level similarity for summaries and translations.
+  • Machine Translation: BLEU (Papineni et al. 2002) is a precision-based metric that assesses machine-generated text quality by comparing n-grams with reference texts.
+In the experiments, thresholds of 0.15, and 0.20 were established for summarization and machine translation tasks to define the conditions required for task completion at each complexity level of the HPF.
+These thresholds allowed for iterative refinement of HPF prompting strategies.
+4.2 Results on Standard Benchmarks: MMLU, GSM8K, and Humaneval
+The evaluation of HPF effectiveness as shown in Figure 4 spans three standard benchmarks: MMLU, GSM8k, and HumanEval.
+On the MMLU benchmark, which tests general knowledge across multiple domains, all models showed notable improvements over their baseline performance.
+Mistral-Nemo 12B demonstrated the most substantial MMLU enhancement (+21.8%), while Claude 3.5 Sonnet achieved a consistent improvement of 3.5%.
+In mathematical reasoning, assessed through GSM8k, the results revealed a correlation with the model scale.
+Larger models like GPT-4 and Claude 3.5 Sonnet showed modest gains (+4.4% and +1.3% respectively), while smaller models exhibited more variable performance.
+The HumanEval benchmark, which assesses code generation capabilities, revealed the most dramatic improvements across all models.
+Mistral 7B achieved an exception 62.5% improvement in HumanEval scores, followed by Mistral-Nemo 12B with an impressive 51.4% improvement, and Gemma-2 9B with a 50.8% enhancement.
+These findings indicate that HPF improves performance across all benchmarks for most of the LLMs, its impact is particularly pronounced in programming tasks, suggesting that the technique may be especially valuable for enhancing code-related capabilities.
+Table 1 highlights the improved performance of various LLMs on MMLU, with all models showing an HPI index below three.
+This indicates that reasoning over most MMLU samples requires minimal cognitive effort for these models, compared to baseline multi-shot CoT methods (5 shot), which typically require more than five examples and are more cognitively demanding according to HPT.
+Interestingly, while Claude 3.5 Sonnet achieves the highest MMLU accuracy, GPT-4o records the best HPI score, showing that minimal cognitive effort does not necessarily equate to the best performance.
+The enhancement in GSM8k is relatively smaller compared to MMLU, with decreased performances for both Mistral 7B and Gemma 7B.
+The high HPI values for Gemma 7B and Mistral 7B indicate that none of the five prompting strategies in HPF posed significant cognitive challenges for these LLMs, highlighting a limitation of the HPF.
+As shown in Table 2, Claude 3.5 Sonnet achieves a perfect pass@1 of 1.00 with low HPI values, outperforming GPT-4o, which scores 0.95 but has a higher HPI.
+Gemma 7B struggles with the lowest pass@1 of 0.79 and the highest HPI of 3.71, indicating a need for more complex prompting strategy.
+Interestingly, HPF significantly enhanced the performance of most LLMs across three benchmark datasets, even when the HPI difference was less than 1 relative to the best performing LLMs.
+This highlights that tailoring the prompting strategy to align with the complexity of each dataset instance can lead to substantial improvements, achieving performance levels comparable to state-of-the-art LLMs such as GPT-4o and Claude 3.5 Sonnet on these benchmarks.
+Figure 4: Performance Comparison of HPT-based Evaluation vs.
+Standard Evaluation: Performance improvements (in %) when using HPT-based evaluation compared to standard evaluation across three benchmarks: MMLU, GSM8k, and HumanEval.
+Positive values indicate performance gains with HPT, while negative values indicate performance decreases.
+The baseline standard evaluation scores are sourced from Hugging Face leaderboard and official research reports.
+DATASETS MMLU GSM8k BoolQ CSQA
+Models HPI Accuracy HPI Accuracy HPI Accuracy HPI Accuracy
+GPT-4o 1.81 91.61 1.71 96.43 1.32 96.82 1.65 92.54
+Claude 3.5 Sonnet 1.84 92.16 1.35 97.72 1.20 99.81 2.01 86.15
+Mistral-Nemo 12B 2.45 89.75 3.01 86.80 1.75 99.87 2.06 90.17
+Gemma-2 9B 2.34 87.28 2.17 91.28 1.30 98.28 1.94 88.86
+Llama-3 8B 2.84 82.63 2.34 86.20 1.37 99.30 2.43 84.76
+Gemma 7B 2.93 83.31 6.70 27.88 1.45 99.42 2.50 83.78
+Mistral 7B 2.89 81.45 5.11 46.93 1.41 98.07 2.49 82.06
+Table 1: HPI (lower is better) and accuracy of LLMs across MMLU, GSM8K, BoolQ, and CSQA datasets.
+Blue indicates datasets where the LLM with the best HPI does not achieve the best performance.
+Green indicates the LLM with the best performance over the maximum number of datasets.
+DATASET HumanEval
+Models HPI Pass@1
+GPT-4o 2.25 0.95
+Claude 3.5 Sonnet 1.04 1.00
+Mistral-Nemo 12B 2.07 0.96
+Gemma-2 9B 1.01 0.91
+Llama-3 8B 1.03 1.00
+Gemma 7B 3.71 0.79
+Mistral 7B 1.10 0.93
+Table 2: HPI (lower is better) and Pass@1 of LLMs on the HumanEval dataset.
+Blue indicates datasets where the LLM with the best HPI does not achieve the best performance.
+Green indicates the LLMs with the best performance over the dataset.
+Figure 5: Hierarchy of prompting strategies with LLM-as-a-Judge framework with GPT-4o as the judge.
+4.3 Results on Other Datasets
+Table 1 presents the performance of LLMs on the BoolQ and CSQA datasets.
+Notably, no significant insights emerge from the results, aside from GPT-4o performing unexpectedly poorly, which contrasts with its typical performance.
+With most LLMs achieving near-perfect scores, the BoolQ dataset appears to lack the complexity needed to serve as an effective benchmark for modern LLMs, as they perform exceptionally well even with minimal cognitive prompting strategies.
+This underscores the utility of HPF in evaluating dataset complexities relative to an LLM, offering researchers valuable insights for designing more challenging and robust benchmarks.
+Table 3 presents the performance of LLMs on IWSLT and SamSum datasets at varying thresholds.
+GPT-4o consistently achieved the highest scores across all thresholds, while most models, except Gemma 7B, performed similarly.
+Interestingly, Claude 3.5 Sonnet, which excelled in reasoning tasks, did not perform as strongly in summarization and translation tasks.
+The threshold selection is guided by the observed performance plateau across most LLMs as the threshold increases.
+For a detailed explanation of the threshold selection process, please refer to Appendix B.
+DATASETS IWSLT SamSum
+HPI BLEU HPI ROUGE-L
+Models 0.15 0.20 0.15 0.20 0.15 0.20 0.15 0.20
+GPT-4o 2.66 3.08 0.32 0.32 1.11 1.21 0.30 0.29
+Claude 3.5 Sonnet 4.63 4.87 0.20 0.20 1.25 1.60 0.23 0.23
+Mistral-Nemo 12B 2.87 3.40 0.27 0.27 1.19 1.47 0.23 0.24
+Gemma-2 9B 4.40 4.75 0.21 0.20 1.30 1.86 0.22 0.22
+Llama-3 8B 3.40 3.92 0.24 0.23 1.30 1.72 0.22 0.22
+Gemma 7B 5.39 5.84 0.08 0.09 3.31 5.03 0.11 0.10
+Mistral 7B 3.52 4.14 0.22 0.22 1.26 1.68 0.21 0.22
+Table 3: HPI (lower is better), BLEU score for IWSLT, and ROUGE-L score for SamSum, of LLMs with thresholds.
+4.4 Complexity Levels with LLM-as-a-Judge
+This study evaluated prompting strategies by assessing how GPT-4o, as the LLM judge, replicates the hierarchical complexity levels of these strategies using a systematic scoring approach across tasks.
+Figure 5 shows a consistent hierarchy with less variability than human judges, indicating a strong alignment between LLM and human judgment.
+These results validate the proposed framework and demonstrate the correspondence between human cognitive principles and LLM behavior.
+Figure 6 shows the scoring distribution across the four HPT rules for each strategy.
+Further details related to dataset specifications and scoring method are in Appendix C.
+Figure 6: Scoring distribution for each of the four rules of the HPT for the prompting strategies in the HPF.
+4.5 Parallels with System 1 and System 2 Thinking
+HPF align closely with the principles of System 1 and System 2 thinking from dual-process cognitive theories (Booch et al.
+2021).
+HPT categorizes tasks and HPF structures prompts based on their cognitive complexity, mirroring how humans allocate cognitive resources.
+For tasks with low cognitive demands, HPF employs simple prompts that parallel System 1 thinking.
+These tasks, like fact recall or basic classification, require minimal reasoning, allowing the LLM to respond quickly and efficiently without extensive computation.
+For instance, asking an LLM to ”identify the capital of a country” is analogous to a person retrieving a familiar fact using System 1.
+In contrast, tasks with high cognitive demands involve prompts that guide the LLM through complex reasoning, abstraction, or multi-step problem-solving—analogous to System 2 thinking.
+Examples include generating logical arguments or solving intricate problems, where deliberate and resource-intensive processes are necessary.
+Just as System 2 engages when a problem exceeds the capacity of System 1, higher levels of HPF are invoked for tasks requiring deeper analysis.
+HPF explicitly measures this transition with HPI, assessing the cognitive load required for each task.
+By tailoring prompting strategies to task complexity, HPF optimizes LLM performance, much like humans adaptively switch between System 1 and System 2 based on the situation.
+This parallel highlights how HPT bridges computational strategies with human-like cognitive models, enabling more nuanced task evaluation and resource allocation.
+4.6 Adaptive HPF
+The Adaptive HPF automates the selection of the optimal complexity level in the HPF using a prompt-selector, Llama-3 8B in a zero-shot setting, bypassing iterative steps.
+Figure 7 shows that Adaptive HPF yields higher HPI but lower evaluation scores than the standard HPF.
+This suggests that Adaptive HPF struggles to select the optimal complexity level, likely due to hallucinations by the prompt-selector when choosing the prompting strategy.
+For more results and ablation studies, see Appendix E.
+Figure 7: HPI of datasets for LLMs in Adaptive HPF.
+5 Conclusion
+The HPT provides a strong and efficient approach for assessing LLMs by focusing solely on the cognitive demands of different tasks.
+The results emphasize that the HPF is effective in evaluating diverse datasets, using the most cognitively effective prompting strategies tailored to task complexity, which results in improved LLM performance across multiple datasets.
+This method not only offers an in-depth insight into LLM’s problem-solving abilities but also suggests that dynamically choosing suitable prompting strategies can enhance LLM performance, setting the stage for future improvements in LLM evaluation methods.
+This study paves the way for formulating and designing evaluation methods grounded in human cognitive principles, aligning them with the problem-solving capabilities of LLMs.
+Additionally, it facilitates the development of more robust benchmarks and in-context learning methodologies to effectively assess LLM performance across various tasks.
+6 Ethical Statement
+The $\texttt{HPI}_{Dataset}$ assigned by experts to the datasets: MMLU, GSM8k, Humaneval, BoolQ, CSQA, IWSLT, and SamSum may introduce bias into the comparative analysis.
+This potential bias stems from the subjective nature of expert scoring, which can be influenced by individual experience and perspective.
+Despite this, the datasets utilized in this study are publicly available and widely recognized in the research community, thereby minimizing the risk of unanticipated ethical issues.
+Nevertheless, it is crucial to acknowledge the possibility of scoring bias to ensure transparency and integrity in the analysis.
+Appendix A Human Annotation and Judgement Policy
+A.1 Human Annotation Policy
+$\texttt{HPI}_{Dataset}$ is introduced to penalize the HPI of tasks or samples unsolvable by the LLM, aligning the framework more closely with human cognitive demands and enhancing its comprehensiveness.
+We implemented a rigorous human annotation process to ensure the quality of $\texttt{HPI}_{Dataset}$ scored by human experts for the datasets.
+Human annotators were tasked with calculating the HPI for each sample in a given dataset.
+The HPI quantifies the cognitive demands imposed on human expert proficiency in completing a task, based on the HPT, where higher values indicate greater cognitive demands.
+Each sample was scored on a scale from 1 (lowest complexity level) to 5 (highest complexity level) for the following criteria:
+  1. 1.
+    Basic Understanding and Reproduction: This criterion evaluates the annotator’s ability to comprehend and accurately reproduce the content.
+  2. 2.
+    Understanding and Interpretation: This criterion assesses the annotator’s depth of understanding and the ability to interpret the information correctly.
+  3. 3.
+    Analysis and Reasoning: This criterion measures the annotator’s ability to analyze the information and apply logical reasoning.
+  4. 4.
+    Application of Knowledge and Execution: This criterion evaluates the annotator’s practical application of knowledge and the execution of tasks based on the relevant knowledge.
+Higher scores for the four rules signify a stronger influence of the respective rules, indicating that completing the task requires greater cognitive effort.
+The $\texttt{HPI}_{Dataset}$ for each dataset, as shown in Table 4, was calculated by taking the mean of the values from these four criteria, acknowledging the challenge of estimating or computing the individual weights of the influence of each rule.
+The Representative Set Size in Table 4 refers to the subset of the dataset evaluated by human annotators, ensuring that the assessment reflects the overall task.
+Human annotation, while time-consuming and costly, provides a gold standard for calibrating the evaluation process of this paper.
+Selecting 5% of the dataset as the representative set size balances quality assessment and feasibility, capturing the dataset’s diversity and ensuring that human annotations encompass a broad range of cases without needing to annotate every sample.
+Dataset Evaluation Set Size Representative Set Size $\texttt{HPI}_{Dataset}$
+MMLU 14500 725 3.03
+GSM8k 1320 66 2.14
+Humaneval 160 8 4.68
 BoolQ 3270 162 1.71
@@
 SamSum 819 40 2.23
-Table 1: HP-Scores of datasets by human experts.
-3.4.2 HP-Score for Manual HPF
-HPF categorizes prompting strategies based on their complexity into five levels: Role Prompting (level 1), Zero-CoT (level 2), 3-CoT (level 3), Least-to-Most Prompting (level 4), and GKP (level 5).
-Each level has a score of the same value as the level, i.e.
-for level x, score = x.
-A LLM’s superior task-solving ability is indicated by a lower HP-Score.
-If the LLM fails to address the task at all levels of prompts, it is assigned a penalty corresponding to its dataset HP-Score.
-If the LLM was able to address the task at level x:
-$\text{hp}_{\text{j}}=x$ (3)
-If LLM was unable to address the task after all m levels:
-$\text{hp}_{\text{j}}=m+\text{HP-Score}_{\text{Dataset}}$ (4)
-HP-Score for Manual HPF is:
+Table 4: $\texttt{HPI}_{Dataset}$ scores across datasets evaluated by human annotators.
+The table lists the evaluation set size, representative set size, and $\texttt{HPI}_{Dataset}$ for various datasets.
+$\texttt{HPI}_{Dataset}$ scores provide a measure of task complexity relative to human annotators.
+A.2 Human Judgement Policy
+To populate the HPF with relevant prompting strategies across a wide range of strategies, human annotators who adhered to the annotation policy for assessing $\texttt{HPI}_{Dataset}$ were instructed to follow a judgment policy for a predefined set of prompting strategies.
+They were instructed to evaluate the influence of the four rules of the HPT on solving the annotated tasks using each prompting strategy, rating their influence as High (H), Moderate (M), or Low (L).
+It’s important to note that a high rating on rule 4 has a greater influence than a high rating on rule 3, and similarly for the other two rules.
+Considering the rating as shown in Table 5 and varying influences of these rules, five prompting strategies that prioritize comprehensive coverage of cognitive demands while ensuring the set optimally widens the variation across complexity levels were selected for populating the HPF.
+Prompting Strategy Rule 1 Rule 2 Rule 3 Rule 4
+Role Prompting L L L L
+Emotion Prompting L L M L
+Zero-shot CoT L L M L
+Meta Prompting M H M L
+Three-shot CoT H H M L
+Five-shot CoT H H H L
+Chain-of-Verification H H H H
+Least-to-Most Prompting H H H L
+Self-Consistency H H H M
+GKP L H H H
+Table 5: Human judgment of influence of the rules of taxonomy on different prompting strategies in solving the tasks of the representative set.
+The ratings are provided based on a voting system involving all human annotators.
+Green represents the prompting strategies selected for populating the complexity levels of the HPF.
+Appendix B Threshold Selection for Summarization and Translation Tasks
+In addition to the 0.15 and 0.20 thresholds presented in the main paper, extended evaluations were conducted on the IWSLT and SamSum datasets using thresholds of 0.25 and 0.30 with GPT-4o, Mistral-Nemo 12B, and Llama-3 to assess the impact of varying thresholds on LLM performance.
+B.1 Summarization Task
+In the summarization task, increasing the threshold evaluates an LLM’s ability to condense content while retaining key information.
+Higher thresholds like 0.25 and 0.30 reveal the trade-offs between conciseness and informativeness.
+However, as shown in Figure 8, there was no significant improvement in ROUGE-L, except for a slight increase with GPT-4o.
+The experiments showed a sharp rise in HPI, reflecting the increased task complexity.
+These results suggest that LLM performance has plateaued, with no further gains at higher thresholds.
+This validates the use of 0.15 and 0.20 thresholds in the main paper are sufficient for optimal LLM performance.
+Figure 8: Comparison of HPI and ROUGE-L score across different threshold values in summarization task.
+B.2 Translation Task
+In machine translation, higher thresholds (0.25 and 0.30) impose stricter evaluations, assessing how well models capture the nuances of the source text.
+Lower thresholds (0.15 and 0.20) focus on general adequacy, while higher ones test performance under more challenging conditions.
+As shown in Figure 9, no BLEU improvements were observed across any LLMs, with models either reaching saturation or showing decreased performance alongside a rapid rise in HPI.
+This validates the selection of 0.15 and 0.20 thresholds in the main paper as sufficient for optimal LLM performance.
+Figure 9: Comparison of HPI and BLEU score across different threshold values in the translation task.
+Appendix C LLM-as-a-Judge
+C.1 Scoring Prompt Template
+The system prompt is designed to guide the LLM judge in evaluating different prompting strategies based on four specific criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Execution.
+Each criterion is scored on a scale of 1-5.
+The evaluation uses GPT-4o as a judge, with the following system prompt:
+You are a judge evaluating different prompting strategies and you need to score these prompting strategies based on pre-defined criteria.
+Different prompting strategies leverage varied amounts of intelligence from the model to achieve the required answer.
+So, assign the scores very carefully based on your analysis of the prompt and its effect on your intelligence to achieve the given answer as well as the number of multi-step prompts which increases the complexity of execution.
+score1: Basic Recall and Reproduction
+Definition: The need of the model to remember and reproduce factual information without interpretation or analysis to answer the prompt
+Range: 1-5
+score2: Understanding and Interpretation
+Definition: The need of the model to comprehend and explain the meaning of information, summarizing or clarifying content to answer the prompt
+Range: 1-5
+score3: Analysis and Reasoning
+Definition: The need for the model to break down complex information, understand relationships, and solve problems using logical reasoning to answer the prompt
+Range: 1-5
+score4: Application of Knowledge and Execution
+Definition: The need for the model to apply knowledge in practical situations, execute multi-step processes, and solve complex tasks to answer the prompt
+Range: 1-5
+C.2 Hybrid Dataset
+The hybrid dataset is composed of 1106 samples uniformly distributed over seven different task-specific datasets, covering a wide range of language understanding and generation tasks.
+This diversity allows for a comprehensive evaluation of the prompting strategies across various problem types.
+The evaluation uses a hybrid dataset composed of samples from various task-specific datasets and each dataset contributes specific types of tasks:
+  1. 1.
+    MMLU (Massive Multitask Language Understanding)
+  2. 2.
+    HumanEval (Code Generation and Completion)
+  3. 3.
+    GSM8K (Grade School Math 8K)
+  4. 4.
+    BoolQ (Boolean Questions)
+  5. 5.
+    CSQA (Commonsense Question Answering)
+  6. 6.
+    IWSLT (International Workshop on Spoken Language Translation)
+  7. 7.
+    SamSum (Dialogue Summarization)
+C.3 Scoring Method
+For each prompting strategy (Role Prompting, Zero-shot CoT, Three-shot CoT, Least to Most Prompting, Generated Knowledge Prompting), the system:
+  1. 1.
+    Applies the prompting strategy to each sample in the hybrid dataset
+  2. 2.
+    Generates an answer using GPT-4o
+  3. 3.
+    Presents the prompt, generated answer, and correct answer to the LLM judge
+  4. 4.
+    Collects scores for each of the four criteria and the system calculates average scores for each criterion across all tasks and datasets.
+This study ensured that both the human judge and the LLM judge utilized the same scoring methodology to eliminate any potential bias in the comparison.
+Appendix D Limitations
+Human Annotation Constraints: A limitation of this study is the reliance on human evaluation for inducing the $\texttt{HPI}_{Dataset}$ penalty into the HPF.
+While this study assessed 5% of the datasets, expanding coverage would offer a more comprehensive analysis.
+However, due to constraints in human resources for manual annotation, we could not include a larger portion.
+Future work could address this by increasing manpower or automating parts of the evaluation process.
+HPF Optimization: The effectiveness of the HPF heavily relies on the quality of the prompts used at each level of the taxonomy.
+Crafting high-quality prompts that accurately reflect the subtleties of each level demands considerable expertise and repeated refinement.
+This study only investigated a limited set of prompting strategies within the HPF, indicating a need for further research into creating diverse structural frameworks and incorporating additional prompting strategies.
+Zero-shot Prompt Selection: HPF dynamically determines the optimal cognitive complexity level by iterating through the framework’s levels, which leads to increased inference time.
+While this study investigated Adaptive HPF for zero-shot prompt selection, it faced considerable hallucinations.
+Future research should focus on automating HPF using fine-tuning or reinforcement learning-based approaches to select the appropriate complexity level without manual iteration.
+This strategy would optimize inference time and improve overall performance.
+Appendix E Adaptive HPF
+E.1 HPI for Adaptive HPF
+The prompt-selector can dynamically select the most suitable prompting strategy for a given task’s complexity from the HPF’s hierarchy of complexity levels.
+To determine the most effective prompting strategy to complete the task, the prompt-selector was given a maximum number of iterations equivalent to the number of levels in the manual HPF.
+The score for ith iteration is i + x, where x is the complexity level by the prompt-selector.
+If the LLM fails to complete the task after all iterations, it is assigned a penalty, $\texttt{HPI}_{Dataset}$.
+Algorithm 2 Prompt-Selector for Adaptive HPF
 […]
-where $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$, x represents the level of HPF at which task is addressed, $m$ represents the total number of levels in HPF, and $n$ denotes the total number of samples in the evaluation set.
-3.4.3 HP-Score for Adaptive HPF
-The prompt-selector can dynamically select the most suitable prompting strategy for a given task’s complexity from the manual HPF’s collection of prompting strategies.
-In order to determine the most effective prompting strategy to address the task, the prompt-selector was given a maximum number of iterations equivalent to the number of levels in the manual HPF.
-The score for ith iteration is i + x, where x is the level of the prompting strategy selected by the prompt-selector.
-If the LLM fails to address the task after all iterations, it is assigned a penalty corresponding to it’s dataset HP-Score.
-If LLM was able to address the task at ith iteration:
-$\text{hp}_{\text{j}}=x+i$ (6)
-If LLM was unable to address the task after all m iterations:
-$\text{hp}_{\text{j}}=m+\text{HP-Score}_{\text{Dataset}}$ (7)
-HP-Score for Adaptive HPF is:
+for sample $j$ in the evaluation dataset do
 […]
-where $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$, x represents the level of HPF selected by prompt-selector at $i$th iteration at which task is addressed, $m$ represents the total number of levels in HPF, and $n$ denotes the total number of samples in the evaluation set.
-4 Results
-4.1 Results for Manual HPF
-Table 2 presents the HP-Scores for the four LLMs: Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B, across previously mentioned datasets with varying threshold values (0.15 and 0.20) for BLEU Papineni et al.
-(2002) and ROUGE Lin (2004).
-The BLEU and ROUGE thresholds at every prompting level serve as the task-solving criteria for each sample of IWSLT and SamSum respectively.
-The complexity at which the LLMs are able to handle the dataset effectively is reflected in the HP-Scores.
-Figure 3: HP-Scores of all datasets and threshold values for LLMs in Manual HPF.
+     for iteration $i=1$ to $m$ do
+         Select prompting strategy at level $x$
+         if LLM completes the task at iteration $i$ then
+[…]
+[…]
+              break
+         end if
+     end for
+     if solved = False then
+[…]
+     end if
+end for
+$\texttt{HPI}_{\texttt{Adaptive}}=\frac{1}{n}\sum_{j=1}^{n}\texttt{HPI\_List}[j]$
+$x$ represents the level of the HPF selected by prompt-selector at $i$th iteration at which the task is addressed, $m$ represents the total number of levels in the HPF, and $n$ denotes the total number of samples in the evaluation set.
 Model BoolQ CSQA IWSLT (0.15) IWSLT (0.20) SamSum (0.15) SamSum (0.20)
-Llama 3 8B 1.3692 2.4281 3.3955 3.9246 1.3018 1.7166
-Mistral 7B 1.4063 2.4929 3.5183 4.1455 1.2590 1.6770
-Phi 3 3.8B 1.3965 2.1714 3.8479 4.4401 2.0925 3.4010
-Gemma 7B 1.4558 2.5053 5.3929 5.8369 3.3137 5.0316
-Table 2: HP-Scores of datasets (with thresholds) across all LLMs for Manual HPF.
-Dataset Metric Threshold Llama 3 8B Phi 3 3.8B Mistral 7B Gemma 7B
-BoolQ Accuracy - 0.99296 0.98685 0.98073 0.99419
-CSQA Accuracy - 0.84766 0.88452 0.82063 0.83783
-IWSLT BLEU 0.15 0.23539 0.19327 0.21629 0.08049
-0.2 0.23091 0.19098 0.21569 0.06807
-SamSum ROUGE-1 0.15 0.29395 0.21389 0.28529 0.16067
-0.2 0.29996 0.22109 0.29365 0.14063
-Table 3: Evaluation scores for datasets across all LLMs for Manual HPF.
-Figure 3 shows that Llama 3 8B achieves lower HP-Scores than other LLMs across all datasets, indicating superior performance and the ability to solve tasks with simpler prompting strategies.
-All four LLMs were better at solving BoolQ than humans, with lower HP-Scores than BoolQ HP-Score by human experts.
-IWSLT was the task where all four LLMs were incapable of solving even at threshold values of 0.15 and 0.20, indicating that LLMs are incapable in solving domain-specific tasks compared to human experts in the respective domains.
-Table 3 shows the evaluation scores for each LLM across all datasets using the Manual HPF.
-Llama 3 8B and Gemma 7B outperform Phi 3 3.8B and Mistral 7B on BoolQ, with all four LLMs achieving near-perfect accuracy.
-Phi 3 3.8B outperformed all other LLMs in solving CSQA.
-Despite having the fewest parameters, phi 3 3.8B achieved the lowest HP-Score among all four LLMs while also achieving the highest accuracy.
-Llama 3 8B has the highest BLEU scores and ROUGE scores across both threshold values for IWSLT and SamSum datasets respectively.
-While Gemma 7B has the lowest performance and the highest HP-Scores for the IWSLT and SamSum datasets.
-Llama 3 8B and Mistral 7B are the most capable and genearlized LLMs, achieving superior performance across all datasets.
-Phi 3 3.8B, even with the fewest parameters, was able to compete with Llama 3 8B and Mistral 7B.
-Gemma 7B even though was able to solve BoolQ and CSQA with superior performance showing it’s reasoning abilities but was unable to achieve good performance across domain-specific datasets like IWSLT and SamSum.
-Model BoolQ CSQA IWSLT (0.15) IWSLT (0.20) SamSum (0.15) SamSum (0.20)
-Llama3 5.2173 5.9136 6.2006 6.2841 5.0316 5.5756
+Llama-3 8B 5.2173 5.9136 6.2006 6.2841 5.0316 5.5756
 Mistra 7B 5.0483 5.9073 6.2478 6.4604 4.7423 5.1336
-Phi 3 3.8B 5.1386 5.6793 6.3955 6.4936 5.0961 5.7778
+Phi-3 3.8B 5.1386 5.6793 6.3955 6.4936 5.0961 5.7778
 Gemma 7B 5.1514 5.5771 6.5947 6.6605 5.7229 6.4347
-Table 4: HP-Scores of datasets (with thresholds) across all LLMs for Adaptive HPF.
-Dataset Metric Threshold Llama3 8B Phi 3 3.8B Mistral 7B Gemma 7B
+Table 6: HPI (lower is better) of LLMs across datasets (with thresholds) for Adaptive HPF.
+Dataset Metric Threshold Llama-3 8B Phi-3 3.8B Mistral 7B Gemma 7B
 BoolQ Accuracy - 0.88577 0.91115 0.91752 0.91166
@@
 0.2 0.24981 0.21580 0.28335 0.16179
-Table 5: Evaluation scores for datasets across all LLMs for Adaptive HPF.
-4.2 Results for Adaptive HPF
-The Adaptive HPF automates the selection of appropriate prompting strategy for each task using a prompt-selector utilizing Llama 3 8B for the experiments.
-Table 4 shows the HP-Scores and Table 5 presents the evaluation scores for the LLMs using the adaptive HPF.
-Figure 4: HP-Scores of all datasets and threshold values for LLMs in Adaptive HPF.
-The findings presented in Figure 4 show that, on all datasets, the adaptive HPF produces HP-Scores for the LLMs that are higher than those obtained with the manual HPF, with decreased evaluation scores.
-Higher HP-Scores demonstrate that adaptive HPF is obviously unable to complete a variety of tasks, with the primary cause being hallucinations when choosing the appropriate level of prompting by the prompt-selector.
-Hallucination of prompt-selector leads to repeatedly choosing irrelevant prompting levels or choosing levels out of the range of manual HPF for several iterations of adaptive HPF, resulting in high HP-Scores.
-This demonstrates that LLMs are incapable of judging their evaluations of various tasks.
-Figure 5: Accuracy scores of BoolQ and CSQA datasets across all LLMs for manual HPF and Adaptive HPF.
-Figure 6: BLEU scores of IWSLT dataset across all LLMs for manual HPF and Adaptive HPF..
-Figure 7: ROUGE-1 scores of SamSum dataset across all LLMs for manual HPF and Adaptive HPF..
-4.3 Comparison between Manual HPF and Adaptive HPF
-Figure 5 illustrates that the manual HPF outperforms the adaptive HPF on the BoolQ dataset across all LLMs.
-The decrease in accuracy with the adaptive HPF demonstrates its inability to select the most appropriate prompting level within its iteration limit.
-Figure 5 shows a significant decline in performance for LLMs in solving CSQA, suggesting that manual HPF is more dependable than adaptive HPF for solving CSQA.
-Despite solving the IWSLT with slightly worse performance as shown in Figure 6 when using the adaptive HPF, LLMs are unable to identify the optimal prompting level indicated by their high HP-Scores.
-Figure 7 shows that both frameworks perform similarly in solving SamSum, with the adaptive HPF showing higher HP-Scores across all LLMs.
-Table 1 shows that the HP-Scores of both manual HPF and adaptive HPF indicate that manual HPF more closely matches human estimation of task complexity, while the adaptive HPF faces difficulty in solving these tasks when compared to both manual HPF and humans.
-Both frameworks have higher HP-Scores when solving IWSLT than domain experts’ HP-Scores, indicating that pre-trained LLMs with more domain-specific knowledge are needed to demonstrate better performance across domain-specific tasks with less complex prompting strategies, reducing the need for these LLMs to be fine-tuned.
-Another important finding from the IWSLT and SamSum evaluation scores is that all four LLMs were unable to perform above a particular evaluation score at higher threshold values, as evidenced by the saturation of their performance across all threshold values.
-Overall, our experiments demonstrate that different tasks don’t always require sophisticated prompting strategies from LLMs; it largely depends on the LLM’s aptitude for the task.
-The HP-Score provides a universal measure of task complexity relative to different agents, offering insights into the capabilities of both LLMs and humans.
-When comparing Manual HPF and Adaptive HPF, the Manual HPF is more reliable and consistently performs well on datasets with lower HP-Scores.
-In contrast, the Adaptive HPF is computationally inefficient due to hallucinations when selecting the prompting level in least number of iterations.
-For other tasks and LLMs, Adaptive HPF requires finetuning of the prompt selector or other techniques to minimize hallucinations and ensure consistent outcomes.
-5 Conclusion
-The HPT, combined with the HPFs, offers a robust and efficient method for evaluating LLMs based entirely on task complexity.
-These findings highlight the effectiveness of the HPF in evaluating the task with the most effective prompting strategies based on task complexity, leading to enhanced LLM performance across various datasets and tasks.
-This approach not only provides a detailed understanding of LLMs problem-solving capabilities but also introduces Adaptive HPF by dynamically selecting the most appropriate prompting strategy, paving the way for future advancements in LLM evaluation techniques.
-6 Limitations
-Restricted Model Evaluation: The study focused on four specific instruction-tuned LLMs: Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B.
-While these LLMs provide valuable insights, the inclusion of a more diverse set of LLMs, both in terms of architecture and scale, would help assess the applicability of the HPT across a broader spectrum of LLMs.
-Restricted Dataset Evaluation: The evaluation of both HPF and adaptive HPF is conducted on four distinct datasets encompassing tasks such as question-answering, reading comprehension, translation, and summarization.
-Incorporating a wider variety of datasets that cover additional tasks would aid in evaluating the HPT’s applicability across a more extensive range of tasks.
-Manual framework optimization: The effectiveness of the HPF heavily relies on the quality of the prompts used at each level of the taxonomy.
-Crafting high-quality prompts that accurately reflect the subtleties of each level demands considerable expertise and repeated refinement.
-This study only investigated a limited set of prompting strategies within the HPF, indicating a need for further research into creating diverse structural frameworks and incorporating additional prompting strategies.
-Adaptive framework optimization: The Adaptive HPF automates the selection of suitable prompting strategies according to task complexity.
-However, the current system depends on a prompt-selector that utilizes an LLM.
-There were numerous hallucinations when choosing the appropriate prompting strategy for the task.
-Further investigation is required to minimize hallucinations in the prompt-selector and to explore alternative methods, such as ensemble techniques or meta-learning approaches, to enhance the efficiency and effectiveness of the adaptive framework.
-7 Ethical Consideration
-The HP-Scores assigned by experts in the domains related to the datasets—BoolQ, CSQA, IWSLT and SamSum—may introduce bias into the comparative analysis.
-This potential bias stems from the subjective nature of expert scoring, which can be influenced by individual experience and perspective.
-Despite this, the datasets utilized in this study are publicly available and widely recognized in the research community, thereby minimizing the risk of unanticipated ethical issues.
-Nevertheless, it is crucial to acknowledge the possibility of scoring bias to ensure transparency and integrity in the analysis.
-Appendix A Large Language Models Used for Evaluation
-The HPF supports multiple LLMs and includes mechanisms for optimizing performance through advanced quantization techniques.
-The experiments were conducted on the following instruction-tuned LLMs and the model description and licenses are discussed in Table 6.
+Table 7: Performance scores of LLMs across datasetsfor Adaptive HPF.
+E.2 Hallucination in Adaptive HPF
+Hallucinations in prompt-selector refer to instances where the LLM generates incorrect or misleading prompting levels or nonsensical information that is not supported by the HPF.
+These hallucinations can occur across various tasks, including question answering, multiple-choice questions, translation, and summarization.
+For the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations (represented by ’…’).
+However, by the fourth iteration, Llama-3 8B manages to answer correctly at Level 2.
+For the CSQA task, prompt-selector exhibits hallucinations initially, shown by Level 4 and Level 0 (not included in HPF) responses.
+Eventually, it corrects itself by the third iteration, providing the correct answer at Level 2.
+For the IWSLT task, prompt-selector demonstrates a consistent pattern of hallucinations across multiple iterations.
+Even though Llama-3 8B attempts the translation at Level 2 multiple times, it ultimately fails to provide a correct translation, indicating a persistent hallucination.
+For the SamSum task, prompt-selector shows initial hallucinations in the first three iterations (Level 4).
+However, by the fourth and fifth iterations, the prompt-selector starts producing lower levels.
+Finally, Llama-3 8B achieves the correct answer at Level 2 in the last iteration .
+The results in Table 6 and Table 7 indicate that the prompt-selector exhibits hallucinations in selecting complexity levels across various tasks and iterations resulting in higher HPI for Adaptive HPF, with performance varying significantly.
+While the LLM can eventually produce correct answers, as seen in the BoolQ and SamSum tasks, it often requires multiple attempts and may still fail in tasks like IWSLT translation.
+E.3 Prompt Template for Prompt-Selector
+The prompt-selector in adaptive HPF selects the prompting level based on the task complexity to address the task.
+Llama-3 8B serves as the prompt-selector in the experiments.
+The prompt template was meticulously designed to ensure maximum clarity, aiming to reduce hallucinations and select the most effective prompting strategy.
+Prompt Template: Choose the most effective prompting strategy among five available strategies for the task.
+Begin with the lowest indexed strategy and progress to higher indexed strategies if the earlier ones are not effective.
+For a given task, the prompting strategies are:
+  • Role Prompting: Defines a role for the model in solving the task.
+  • Zero-shot Chain of Thought prompting: Stimulates reasoning and problem-solving by including the phrase ’Let’s think step by step’ without offering previous examples related to the task.
+  • Three-shot Chain of Thought prompting: Offers three examples related to the task to guide the model’s reasoning process.
+  • Least-to-most prompting: Uses a sequential method to derive essential insights from the task to solve it.
+  • Generated Knowledge Prompting: Integration and application of external knowledge to accomplish the task. The external knowledge is generated using some other model based on the task.
+Select only the index (do not provide the name) of the most effective prompting strategy.
+Appendix F Computational Budget
+All evaluation experiments and ablation studies were conducted on V100 GPUs (16GB and 32GB variants), utilizing a total of around 9,000 computation hours, this equates to approximately 1.125 petaflop-hours of computational resources.
+Appendix G Large Language Models Used for Evaluation
+The HPF supports leading open source and proprietary LLMs and includes mechanisms for optimizing performance through advanced quantization techniques.
+The experiments were conducted on the following instruction-tuned LLMs, and the model description and licenses are discussed in Table 8.
 Model License Type Usage Restrictions
-Llama 3 8B Research License Non-commercial, requires attribution, ethical guidelines
-Phi 3 3.8B Open-source Broad use allowed, must include original license and notices
+GPT-4o Proprietary Commercial use requires paid API access, subject to OpenAI’s terms of service
+Claude 3.5 Sonnet Proprietary Commercial use requires paid API access, subject to Anthropic’s terms of service
+Mistral-Nemo 12B Proprietary Usage likely restricted to authorized partners or specific use cases
+Gemma-2 9B Research License Non-commercial use only, research purposes
+Llama-3 8B Research License Specific restrictions may apply, typically for non-commercial research use
 Mistral 7B Open-source Broad use allowed, must include original license and notices
-Gemma 7B Open-source/Research Depending on license, may have non-commercial restrictions or broad use allowed
+Gemma 7B Open-source/Research Depending on the license, may have non-commercial restrictions or broad use allowed
+Phi-3 3.8B Open-source Broad use allowed, must include original license and notices
 Table 8: License information for LLMs used in the experiments.
-The selection of Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B offers a comprehensive evaluation set that spans different model sizes, architectures, and use cases, ensuring the HPF framework is robust and generalizable.
-Llama 3 8B excels in language understanding and generative tasks, achieving top-tier performance across benchmarks such as MMLU, HumanEval, GSM-8K, BoolQ, and CommonSenseQA.
-This model’s proficiency in diverse tasks, including complex mathematical reasoning and commonsense reasoning, underscores its high-quality text generation capabilities.
-Phi 3 3.8B, optimized for environments with limited computational resources, performs competitively on benchmarks like GSM-8K and BoolQ, making it ideal for real-time applications due to its efficiency and quick response times.
-Mistral 7B balances parameter count and performance, excelling in MMLU, CommonSenseQA, and HumanEval, showcasing versatility in language understanding and code-related tasks.
-Gemma 7B, specialized for particular domains, excels in niche areas like medical text understanding and technical documentation.
-While not always the top performer in general benchmarks, Gemma 7B’s specialization complements the general-purpose nature of the other LLMs, ensuring the HPF framework is applicable to a wide range of scenarios and tasks.
-By including these LLMs, the effectiveness of the HPF is highlighted, demonstrating its ability to accommodate a wide range of scenarios and tasks, from general-purpose language understanding and generation to specialized domain-specific applications.
-This comprehensive evaluation ensures that HPF is a robust and generalizable framework in the landscape of LLM evaluation and prompting frameworks.
-The quantization configuration for each of the instruction-tuned LLM involves reducing the precision of the LLM’s weights and computations, allowing for a more compact representation and faster processing.
-The LLMs were loaded in a 4-bit precision format which significantly reduces the memory footprint without compromising accuracy.
-By using half-precision floating point computations (float16), the LLMs maintain high computational efficiency.
-Additionally, the quantization type employed is NF4, which provides a balanced approach between model size and performance.
-The use of double quantization further refines this process, ensuring that the LLMs operate efficiently even on resource-constrained hardware.
-The generation configuration is carefully designed to optimize the quality of text output during evaluation.
-This involves setting parameters that influence the behavior of the LLMs during the text generation process.
-For instance, the maximum number of tokens that can be generated in a single run is capped (1024 tokens) to ensure concise and relevant outputs.
-The temperature parameter was set to 0.6 to control the randomness of the predictions; lower temperature results in more deterministic and focused outputs.
-Top-p (p=0.9) sampling allows the LLMs to explore various possible continuations for a given input.
-Finally, the repetition penalty is applied to discourage the LLMs from repeating the same phrases or sentences, thus ensuring more coherent and varied text generation.
-Appendix B Prompt Templates
-B.1 Level 1: Role Prompting
-Role prompting is the most basic level of interaction with a LLM, where the LLM is given a specific role or task to perform without additional context or examples.
-This technique relies solely on the initial instruction to guide the LLM’s response.
-For instance, if the LLM is asked to "act as a translator", it will attempt to translate text based on its training data.
-This method is straightforward but may lack depth and context, leading to less accurate or nuanced outputs.
-Table 7 presents the prompt templates utilized for Role prompting across all datasets in our experiments.
+The LLMs were loaded in 4-bit precision format, with a maximum generation limit of 1024 tokens per run to ensure concise outputs.
+The temperature was set to 0.6 to control prediction randomness, while top-p sampling (p=0.9) enabled the exploration of diverse continuations.
+Additionally, a repetition penalty was applied to discourage the generation of repeated phrases, promoting coherent and varied text output.
+Appendix H Prompt Templates
+H.1 Level 1: Role Prompting
+Role prompting represents the most basic interaction with an LLM, assigning it a specific role or task without additional context or examples.
+This approach relies solely on the initial instruction to guide responses.
+For instance, asking the LLM to “act as a translator” prompts it to translate text based on its training data.
+While straightforward, this method may lack depth, resulting in less accurate or nuanced outputs.
+Table 9 shows the prompt templates used for role prompting across all datasets in the experiments.
 Dataset Prompt
@@
 SamSum Summarize the Dialogue: ‘‘dialogue" as a Summarizer.
+GSM8k Based on the question: "question", calculate the numerical answer to the question as an expert mathematician.
+HumanEval Complete the given code based on the mentioned constraints: "code" as an expert programmer.
+MMLU Choose the answer: ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4" as a critical thinker.
 Table 9: Prompt templates of different datasets for Role Prompting.
-B.2 Level 2: Zero-shot Chain-of-Thought Prompting
-Zero-shot Chain-of-Thought prompting builds upon basic role prompting by asking the LLM to generate a reasoning process for a given task, even though it has not been explicitly trained on similar examples.
-In this approach, the LLM attempts to break down the problem and solve it step-by-step using its internal knowledge base.
-This technique leverages the LLM’s ability to infer and reason, improving the quality of the response by encouraging logical progression and coherence in the thought process.
-Table 8 presents the prompt templates utilized for Zero-CoT across all datasets in our experiments.
+H.2 Level 2: Zero-shot Chain-of-Thought Prompting
+Zero-shot Chain-of-Thought (CoT) prompting enhances basic role prompting by requiring the LLM to generate a reasoning process for a task, despite not being explicitly trained on similar examples.
+This method encourages the LLM to break down the problem and solve it step-by-step using its internal knowledge, improving response quality through logical progression and coherence.
+Table 10 displays the prompt templates used for Zero-CoT across all datasets in the experiments.
 Dataset Prompt
@@
 SamSum Summarize the Dialogue: ‘‘dialogue". Let’s think step by step.
+GSM8k Based on the question: "question", calculate the numerical answer to the question. Let’s think step by step.
+HumanEval Complete the given code based on the mentioned constraints: "code". Let’s think step by step.
+MMLU Choose the answer: ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4". Let’s think step by step.
 Table 10: Prompt templates of different datasets for Zero-shot Chain-of-Thought Prompting.
-B.3 Level 3: Three-Shot Chain-of-Thought Prompting
-Three-shot Chain-of-Thought prompting advances the zero-shot approach by providing the LLM with three examples of the task at hand, including the reasoning steps taken to arrive at the solution.
-These examples serve as a guide, helping the LLM understand the structure and logic required to solve similar problems.
-By seeing how previous tasks were approached and resolved, the LLM can better mimic the process, leading to more accurate and contextually appropriate responses.
-Table 9 presents the prompt templates utilized for 3-CoT across all datasets in our experiments.
+H.3 Level 3: Three-Shot Chain-of-Thought Prompting
+Three-shot Chain-of-Thought (CoT) prompting builds on the zero-shot approach by providing the LLM with three task examples, including the reasoning steps used to reach the solution.
+These examples help the LLM grasp the required structure and logic, enabling it to better replicate the problem-solving process and produce more accurate, contextually relevant responses.
+Table 11 shows the prompt templates used for 3-CoT across all datasets in the experiments.
 Dataset Prompt
@@
 Summarize the Dialogue: "dialogue".
-Table 9: Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting
-B.4 Level 4: Least-to-Most Prompting
-Least-to-most prompting is a sophisticated technique that involves gradually increasing the complexity of prompts presented to the LLM.
-The process begins with simpler, more manageable tasks and progressively moves to more complex challenges.
-This method helps the LLM build confidence and competence in solving problems, as it can leverage the insights gained from simpler prompts to tackle more difficult ones.
-It enhances the LLM’s ability to generalize from straightforward examples to more intricate scenarios.
-Table 10 presents the prompt templates utilized for Least-to-Most Prompting across all datasets in our experiments.
+GSM8k Based on the question:"gsm8k_question1", calculate the numerical answer to the question.Answer: "gsm8k_ans1".Based on the question:"gsm8k_question2", calculate the numerical answer to the question.Answer: "gsm8k_ans2".
+Based on the question:"gsm8k_question3", calculate the numerical answer to the question.Answer: "gsm8k_ans3".
+Based on the question: "question", calculate the numerical answer to the question.
+HumanEval Complete the given code based on the mentioned constraints: "humaneval_code1", Code: "humaneval_sol1".
+Complete the given code based on the mentioned constraints: "humaneval_code2", Code: "humaneval_sol1".Complete the given "code" based on the mentioned constraints: "humaneval_code3", Code: "humaneval_sol3".
+MMLU Choose the answer for the question: "mmlu_ques1" A.
+[AND, NOT] B.
+[NOT, OR] C.
+[AND, OR] D.
+[NAND] Answer: C.
+Explanation: "mmlu_exp1".
+Choose the answer for the question "mmlu_ques2" A.
+The defendant’s statement was involuntary.
+B.
+The defendant’s statement was voluntary.
+C.
+The defendant was not in custody when the statement was made.
+D.
+The statement was not made in response to a known police interrogation.
+Answer: A, Explanation: "mmlu_exp2".
+Choose the answer for the question: "mmlu_ques3" .
+A.
+Wrong, Wrong.
+B.
+Wrong, Not wrong C.
+Not wrong, Wrong D.
+Not wrong, Not wrong.
+Answer: B Explanation: "mmlu_exp3".
+Choose the answer."question" ‘‘question",A.
+‘‘option 1",B.
+‘‘option 2",C.
+‘‘option 3",D.
+‘‘option 4".
+Table 11: Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting.
+H.4 Level 4: Least-to-Most Prompting
+Least-to-most prompting is an advanced technique that gradually increases prompt complexity, starting with simpler tasks and progressing to more complex challenges.
+This method allows the LLM to build confidence and leverage insights from easier prompts to tackle harder ones, enhancing its ability to generalize from straightforward examples to intricate scenarios.
+Table 12 displays the prompt templates used for Least-to-Most Prompting across all datasets in the experiments.
 Dataset Prompt
@@
 prompt 4: Refine this draft summary to make it more concise and coherent, ensuring it captures the essence of the text: "dialogue".
-Table 10: Prompt templates of different datasets for Least-to-Most Prompting
-B.5 Level 5: Generated Knowledge Prompting
-Generated Knowledge prompting represents one of the most complex prompting techniques present in the HPF, where the LLM not only addresses the given task but also incorporates additional relevant information to enhance its response.
-This method involves prompting another LLM to produce auxiliary knowledge that can be used to better understand and solve the problem.
-By creating a rich context and leveraging self-generated insights, the LLM can provide more detailed, accurate, and nuanced answers, showcasing the full extent of its capabilities.
-Table 11 presents the prompt templates utilized for Least-to-Most Prompting across all datasets in our experiments.
+GSM8k Analyze the question: "question".
+Break the question into sub-problems: "question".
+Calculate answers for the subproblems of the question: "pred".
+Calculate the numerical answer to this question: "question" based on the previous calculations: "pred"
+HumanEval Analyze the code: "code".
+Break the question into sub-problems: "code".
+Complete code for the subproblems of the code: "pred".
+Complete the code based on the mentioned constraints: "code" based on the previous calculations: "pred"
+MMLU Analyze the question: "question".
+Elaborate about each option for the question: "question", options: A.
+"option 1" B.
+"option 2" C.
+"option 3" D.
+"option 4".
+Based on the analysis: "question", Discard wrong answers among the options: A.
+"option 1" B.
+"option 2" C.
+"option 3" D.
+"option 4".
+Table 12: Prompt templates of different datasets for Least-to-Most Prompting.
+H.5 Level 5: Generated Knowledge Prompting
+Generated Knowledge prompting is one of the most complex techniques in HPF, where the LLM not only addresses the task but also integrates relevant additional information to enhance its response.
+This method prompts another LLM to produce auxiliary knowledge, creating a richer context for understanding and solving the problem.
+By leveraging self-generated insights, the LLM can deliver more detailed, accurate, and nuanced answers.
+Table 13 shows the prompt templates used for Generated Knowledge Prompting across all datasets in the experiments.
 Dataset Prompt
@@
 IWSLT inference prompt: Translate "english text": to French using definitions of the keywords:"knowledge"
-knowledge generatiuon prompt: Generate definitions in french of each word in the text: "english text".
+knowledge generation prompt: Generate definitions in french of each word in the text: "english text".
 SamSum inference prompt: Summarize the Dialogue: "dialogue" using the interpretation of the dialogue:"knowledge"
 knowledge generation prompt: Generate interpretation about the dialogue: "dialogue".
-Table 11: Prompt templates of different datasets for Generated Knowledge Prompting
-B.6 Adaptive HPF
-The prompt-selector in adaptive HPF selects the prompting level based on the task complexity to address the task.
-Llama 3 8B serves as the prompt-selector in the experiments.
-The prompt template was meticulously designed to ensure maximum clarity, aiming to reduce hallucinations and select the most effective prompting strategy.
-Prompt Template: Choose the most effective prompting strategy among five available strategies for the task.
-Begin with the lowest indexed strategy and progress to higher indexed strategies if the earlier ones are not effective.
-For a given task, the prompting strategies are:
-  • Role Prompting: Defines a role for the model in solving the task.
-  • Zero-shot Chain of Thought prompting: Stimulates reasoning and problem-solving by including the phrase ’Let’s think step by step’ without offering previous examples related to the task.
-  • Three-shot Chain of Thought prompting: Offers three examples related to the task to guide the model’s reasoning process.
-  • Least-to-most prompting: Uses a sequential method to derive essential insights from the task in order to solve it.
-  • Generated Knowledge Prompting: Integration and application of external knowledge to accomplish the task. The external knowledge is generated using some other model based on the task.
-Select only the index (do not provide the name) of the most effective prompting strategy.
-Figure 8: HP-Scores across all datasets and LLMs for manual HPF and Adaptive HPF.
-Appendix C Human Annotation Policy
-To ensure the quality of HP-Scores scored by human experts for the datasets, we implemented a rigorous human annotation process.
-Human annotators were tasked with calculating the HP-Score for each sample in a given dataset.
-The HP-Score relative to human experts is a measure of the expert capability in addressing the task based on criteria of HPT.
-Each sample was scored on a scale from 1 (Lowest prompting level) to 5 (Highest prompting level) for the following criteria:
-C.1 Criteria
-  1. 1.
-    Basic Understanding and Reproduction
-    • This criterion evaluates the annotator’s ability to comprehend and accurately reproduce the content.
-A high score indicates clear and precise reproduction of the information is required for solving the task.
-  2. 2.
-    Understanding and Interpretation
-    • This criterion assesses the depth of understanding and the ability to interpret the information correctly.
-A high score indicates a detailed and precise understanding of the task is necessary to solve the task.
-  3. 3.
-    Analysis and Reasoning
-    • This criterion measures the ability to analyze the information and apply logical reasoning. A high score demonstrates the need to analyze the information to solve the task.
-  4. 4.
-    Application of Knowledge and Execution
-    • This criterion evaluates the practical application of knowledge and the execution of tasks based on the relevant knowledge.
-A high score indicates effective application and the requirement of knowledge to solve the task.
-C.2 Scoring Process
-Each sample of every dataset was scored individually by the annotators on a scale from 1 to 5, where:
-  • 1 = Poor
-  • 2 = Fair
-  • 3 = Good
-  • 4 = Very Good
-  • 5 = Excellent
-The final HP-Score for each dataset was calculated by taking the mean of the values from these four criteria.
-This approach ensures a comprehensive evaluation of each sample by considering multiple performance criteria.
-Appendix D Additional Results
-D.1 HP-Scores Interpretation
-Figure 8 visualizes HP-Scores for different datasets across various LLMs, comparing manual and adaptive frameworks. Here’s a detailed summary:
-For the BoolQ dataset, the manual HPF HP-Scores are generally lower across all LLMs compared to the Adaptive HPF.
-The adaptive scores are significantly higher, indicating better performance with the manual HPF for solving BoolQ.
-For the CSQA dataset, manual HPF HP-Scores are moderate, slightly higher than those for BoolQ.
-However, similar to BoolQ, the Adaptive HPF HP-Scores are consistently higher, again showing the superiority of the manual HPF over adaptive HPF in solving CSQA.
-In comparison to the other datasets, the manual HPF HP-Scores for the IWSLT (0.15) and IWSLT (0.20) datasets both show a significant increase.
-This increase is particularly noticeable for LLMs such as Phi 3 3.8B and Gemma 7B, which indicate a lack of ability to handle IWSLT.
-With comparable evaluation scores to manual HPF, which suggests no improvement in adaptive HPF, the adaptive HPF HP-Scores continue to be consistently high.
-For both SamSum (0.15) and SamSum (0.20) datasets, the manual HPF HP-Scores are notably high for LLMs such as Phi 3 3.8B and Gemma 7B.
-Adaptive HPF HP-Scores are consistently high compared to manual HPF HP-Scores.
-As evidenced by their lower HP-Scores, manual HPF routinely outperforms adaptive HPF in all tasks and LLMs, including BoolQ, CSQA, and SamSum, where it can outperform human experts.
-With the exception of solving SamSum (0.15) by Mistral 7B in Adaptive HPF, no LLM was able to complete any task with a HP-Score of less than 5, suggesting the need for additional prompting strategies in HPF or improved framework to avoid hallucinations in selecting prompting level by prompt-selector.
-D.2 Hallucination in Adaptive HPF
-Hallucinations in prompt-selector refer to instances where the LLM generates incorrect or misleading prompting levels, or nonsensical information that is not supported by HPF.
-These hallucinations can occur across various tasks, including question answering, multiple-choice questions, translation, and summarization.
-Below, we discuss the occurrence of hallucinations in the Llama 3 8B which was utilized as prompt-selector in our experiments across different tasks and iterations, highlighting how the LLM’s performance varies and where it tends to hallucinate.
-As shown in Figure 9, for the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations (represented by ’…’).
-However, by the fourth iteration, Llama 3 8B manages to answer correctly at Level 2.
-For the CSQA task, prompt-selector exhibits hallucinations initially, shown by Level 4 and Level 0 (not included in HPF) responses.
-Eventually, it corrects itself by the third iteration, providing the correct answer at Level 2.
-For the IWSLT task, prompt-selector demonstrates a consistent pattern of hallucinations across multiple iterations.
-Even though Llama 3 8B attempts the translation at Level 2 multiple times, it ultimately fails to provide a correct translation, indicating a persistent hallucination.
-For the SamSum task, prompt-selector shows initial hallucinations in the first three iterations (Level 4).
-However, by the fourth and fifth iterations, the prompt-selector starts producing lower levels, finally, Llama 3 8B achieving correct answer at Level 2 in the last iteration .
-The results indicate that Llama 3 8B exhibits hallucinations in selecting prompting levels across various tasks and iterations resulting in higher HP-Scores for Adaptive HPF, with performance varying significantly.
-While the LLM can eventually produce correct answers, as seen in the BoolQ and SamSum tasks, it often requires multiple attempts and may still fail in tasks like IWSLT translation.
-Figure 9: Hallucination of prompt-selector in choosing the most effective prompting level to solve corresponding tasks by Llama 3 8B.
-Appendix E Additional Experiments
-E.1 Increasing Threshold for IWSLT and SamSum Datasets
-When comparing the performance of various LLMs, Llama 3-8B stands out for its superior results on the IWSLT and SamSum tasks, consistently performing better in both manual and adaptive frameworks.
-At a threshold of 0.2, the performance of all LLMs reached saturation for these tasks, indicating that increasing the threshold further did not yield improvements.
-Both adaptive and manual frameworks struggled with thresholds higher than 0.2, failing to meet or exceed this benchmark for most samples within their iterations.
-Even the most capable LLM, Llama 3-8B, was unable to address the majority of samples across any prompting level within the HPF.
-This demonstrates the limitations of current HPFs at higher thresholds and highlights the ongoing challenges in achieving better performance in LLMs.
-E.2 Adaptive HPF with Varying Prompt-selectors
-Experiments demonstrated that Adaptive HPF is most effective when Llama 3 8B is used as a prompt-selector compared to LLMs, attributed to its ability to produce the lowest number of hallucinations when selecting the prompting level within the iterations of the adaptive HPF.
-Conversely, Gemma 7B, which exhibited numerous hallucinations across most tasks, was found to be the least effective adaptive HPF.
-Consequently, the choice of the prompt-selector emerges as a crucial factor influencing the efficiency of the adaptive HPF.
+GSM8k Based on the question:"question", calculate the numerical answer to the question using an interpretation of the question:"pred"
+HumanEval Complete the code based on the mentioned constraints:"code" using knowledge of the constraints:"pred"
+MMLU Choose the answer. "question", options: A. "option 1" B. "option 2" C. "option 3" D. "option 4" using knowledge of the question:"pred"
+Table 13: Prompt templates of different datasets for Generated Knowledge Prompting.

### v5

@@
 Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models Aligned with Human Cognitive Principles
-Abstract
+Abstract.
 Assessing the effectiveness of large language models (LLMs) in performing different tasks is crucial for understanding their strengths and weaknesses.
@@
 It assesses the complexity of tasks with the Hierarchical Prompting Index (HPI), which demonstrates the cognitive competencies of LLMs across diverse datasets and offers insights into the cognitive demands that datasets place on different LLMs.
-This approach enables a comprehensive evaluation of an LLM’s problem-solving abilities and the intricacy of a dataset, offering a standardized metric for task complexity.
-Extensive experiments with multiple datasets and LLMs show that HPF enhances LLM performance by 2%-63% compared to baseline performance, with GSM8k being the most cognitively complex task among reasoning and coding tasks with an average HPI of 3.20 confirming the effectiveness of HPT.
-To support future research and reproducibility in this domain, the implementations of HPT and HPF are available here.
-Codehttps://github.com/devichand579/HPT
-1 Introduction
+This approach enables a comprehensive evaluation of LLM’s problem-solving abilities and the intricacy of a dataset, offering a standardized metric for task complexity.
+Extensive experiments with multiple datasets and LLMs show that HPF enhances LLM performance by $2\to 63\%$ compared to baseline performance, with GSM8k being the most cognitively complex task among reasoning and coding tasks with an average HPI of $3.20$ confirming the effectiveness of HPT.
+To support future research in this domain, the implementations of HPT and HPF are publicly available [^1] Code and Experiments.
+Prompting Taxonomy, Cognitive Demands, Prompt Optimization
+isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Cognitive science
+1. Introduction
 Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements in a wide range of applications.
@@
 To achieve a more balanced evaluation framework, it is essential to account for both the task-solving ability of LLMs and the varying cognitive complexities of the dataset samples.
-This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of task complexity.
-Figure 1: Hierarchical Prompting Framework includes five distinct prompting strategies, each designed for different levels of task complexity to ensure the appropriate prompt is selected for the given task.
+This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of sample task complexity.
+[…]
+Hierarchical Prompting Framework diagram.
+Figure 1.
+The Hierarchical Prompting Framework includes five distinct prompting strategies, each designed for different levels of task complexity to ensure the appropriate prompt is selected for the given task.
 A $\checkmark$ indicates task completion, while a $\times$ signifies task incompletion.
-Within this study, complexity is defined as the cognitive demands associated with solving a task or the cognitive load introduced by a prompting strategy on LLMs.
-Henceforth, the term ”complexity” will be applied solely in this context.
-Task complexity, within the realm of human cognition, pertains to the cognitive requirements that a task imposes, which includes the diverse levels of mental effort necessary for processing, analyzing, and synthesizing information.
-According to Sweller (1988), tasks become more complex as they require greater cognitive resources, engaging working memory in more demanding processes such as reasoning and problem-solving.
+This study defines complexity as the cognitive demands imposed by a task or the cognitive load introduced by a prompting strategy on LLMs.
+Task complexity in human cognition reflects the mental effort required for processing, analyzing, and synthesizing information.
+As Sweller (1988) noted, complexity increases with greater cognitive resource demands, engaging working memory in reasoning and problem-solving.
 Similarly, Anderson et al.
-(2014) highlights that human cognitive abilities span a continuum from basic recall to higher-order thinking, with increasing difficulty correlating to tasks that demand analysis, synthesis, and evaluation.
-When applied to LLMs, the complexity of prompting strategies can be systematically evaluated by mapping them onto this human cognitive hierarchy.
-This alignment allows for an assessment of how LLMs perform tasks that reflect varying degrees of cognitive load, thereby providing a structured framework for understanding the cognitive demands associated with various tasks.
-By imposing this cognitive complexity framework on LLMs, this paper establishes a universal evaluation method, grounded in human cognitive principles, that enables more precise comparisons of model performance across tasks with varying levels of difficulty.
+(2014) describes cognitive abilities as a continuum, from basic recall to higher-order thinking, with difficulty rising for tasks requiring analysis, synthesis, and evaluation.
+By mapping LLM prompting strategies onto this hierarchy, we systematically assess how LLMs handle varying cognitive loads.
+This framework provides a structured, cognitively grounded method for evaluating model performance across tasks of differing complexity.
+This study is directed by the following research questions:
+Research Questions: RQ1: Does cognitively inspired strategic selection of prompts enable small language models (SLMs) to match the performance of LLMs?
+RQ2: How can cognitive demand measurements of LLMs guide model selection and deployment decisions beyond traditional metrics?
+RQ3: How can we align prompt complexity with task demands to optimize both computational efficiency and performance?
 This paper introduces the HPT, a set of rules that maps the human cognitive principles for assessing the complexity of different prompting strategies.
 It employs the HPF shown in Figure 1, a prompt selection framework that selects the prompt with the optimal cognitive load on LLM required in solving the task.
-HPF enhances the interaction with LLMs, and improves performance across various tasks by ensuring prompts resonate with human cognitive principles.
-The main contributions of this paper are as follows:
-  1. 1.
-    Hierarchical Prompting Taxonomy (HPT): The paper proposes HPT, a set of rules that maps prompting strategies onto human cognitive principles, thereby enabling a universal measure of task complexity for LLMs.
-  2. 2.
-    Hierarchical Prompting Framework (HPF): The HPF is a framework designed to select the most effective prompt from five distinct prompting strategies, which optimizes the cognitive load on LLMs during task completion.
-This framework not only facilitates a more accurate evaluation of LLMs but also enhances their performance, delivering more transparent insights.
-  3. 3.
+The main contributions of this work are:
+  • Hierarchical Prompting Taxonomy (HPT): The paper introduces HPT, rules mapping prompting strategies to human cognitive principles, enabling a universal measure of LLMs’ task complexity.
+  • Hierarchical Prompting Framework (HPF): The HPF framework selects the best prompt from five strategies to optimize LLMs’ cognitive load, improving evaluation and performance transparency.
   • Hierarchical Prompting Index (HPI): HPI [^2] HPI can be quantitatively assessed to analyze the cognitive abilities of an LLM and the cognitive demands imposed by datasets on LLMs, as both factors are interchangeably related to the complexity of tasks.
-quantitatively evaluates the task complexity of LLMs across various datasets, offering insights into the cognitive demands that each task imposes on different LLMs.
-HPF can be effectively compared to an ”Open Book” examination as shown in Figure 2, where questions represent tasks and textbooks serve as prompting strategies.
-In this analogy, the exam questions vary in complexity, from simple factual recall to intricate analytical problems, analogous to the tasks in HPT that are evaluated based on their cognitive demands.
-Similarly, textbooks provide structured guidance for solving these questions, much like the HPF organizes prompts in increasing levels of complexity to support LLMs.
-For example, a straightforward glossary lookup corresponds to a low-complexity task, while solving a multi-step analytical problem requiring synthesis of concepts represents a high-complexity task.
-The effort a student invests in answering a question mirrors the HPI, which measures the cognitive load placed on the LLM.
-Just as students perform better with structured resources like textbooks, LLMs improve with well-designed hierarchical prompting strategies, enabling them to tackle progressively complex tasks effectively.
-Figure 2: Analogical framework comparing the HPF with ”Open Book” examination methodology.
+quantitatively assesses LLMs’ task complexity across datasets, revealing cognitive demands on various LLMs.
+Figure 2.
+Analogical framework comparing the HPF with "open book" examination methodology.
 The diagram illustrates how HPF components (below) mirror traditional educational assessment elements (above), with parallel relationships between task complexity levels, resource utilization (prompts/textbooks), and performance metrics (HPI/student effort).
-This comparison demonstrates how LLM task complexity scales similarly to educational assessment complexity, from simple lookup tasks to complex synthesis problems
+This comparison demonstrates how LLM task complexity scales similarly to educational assessment complexity, from simple lookup tasks to complex synthesis problems.
+HPF can be compared to an "open book" exam (see Figure 2), with tasks analogous to questions and prompting strategies akin to textbooks.
+The exam questions, ranging from basic recall to complex analysis, parallel the cognitive challenges in HPT tasks.
+Similarly, textbooks offer structured support, much like HPF, which arranges prompts by complexity to assist LLMs.
+A glossary lookup represents a task with low complexity, whereas solving a multi-step analytical problem indicates high complexity.
+The effort exerted by a student is similar to HPI, which measures the cognitive demand on LLMs.
+Just as structured learning materials improve students’ performance, carefully crafted hierarchical prompts help LLMs in addressing increasingly complex tasks more effectively.
 The remainder of the paper is structured as follows: Section 2 reviews the related work on prompting and evaluation in LLMs.
@@
 Section 5 concludes the paper.
+Section 6 discusses the limitations of the work.
 Section 7 discusses the ethical impact of the work.
-2 Related Work
+2. Related Work
 The advent of LLMs has revolutionized NLP by demonstrating significant improvements in few-shot and zero-shot learning capabilities.
@@
 (2021) surveyed prompt-based learning, emphasizing the role of prompt engineering in leveraging pre-trained models for few-shot and zero-shot adaptation to new tasks with minimal labeled data.
-2.1 Prompt Engineering
+2.1. Prompt Engineering
 Prompting plays a vital role in unlocking the full potential of LLMs.
@@
 Recent research has explored various approaches to prompting and reasoning evaluation in LLMs.
-Chain-of-Thought (CoT) prompting (Wei et al.
-2022b) elicits step-by-step reasoning, improving performance on complex tasks.
-Specializing smaller models (Fu et al.
-2023) and using large models as reasoning teachers (Ho, Schmid, and Yun 2022) have demonstrated the potential for enhancing reasoning capabilities.
+Chain-of-Thought (CoT) prompting (Wei et al., 2022b) elicits step-by-step reasoning, improving performance on complex tasks.
+Specializing smaller models (Fu et al., 2023) and using large models as reasoning teachers (Ho et al., 2022) have demonstrated the potential for enhancing reasoning capabilities.
 Emergent abilities in LLMs, which appear suddenly at certain scale thresholds, have also been a topic of interest.
@@
 (2024) introduced role-play prompting to improve zero-shot reasoning by constructing role-immersion interactions, which implicitly trigger chain-of-thought processes and enhance performance across diverse reasoning benchmarks.
-Progressive-hint prompting (Zheng et al.
-2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses.
-Metacognitive prompting (Wang and Zhao 2024) incorporates self-aware evaluations to enhance understanding abilities.
-These works collectively highlight the advancements in leveraging LLMs through innovative prompting techniques, addressing their emergent abilities, reasoning capabilities, interaction strategies, robustness, and evaluation methodologies.
-Despite significant advancements, the current LLM research reveals several limitations, particularly in terms of prompt design, handling complex reasoning tasks, and evaluating model performance across diverse scenarios.
-While promising, the emergent abilities of LLMs often lack predictability and control, and the robustness of these LLMs in the face of misleading prompts remains a concern.
-2.2 Prompt Optimization and Selection
+Progressive-hint prompting (Zheng et al., 2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses.
+Metacognitive prompting (Wang and Zhao, 2024) incorporates self-aware evaluations to enhance understanding abilities.
+These studies highlight progress in using innovative prompting techniques to improve LLMs’ emergent abilities, reasoning, interaction strategies, robustness, and evaluation.
+Yet, challenges persist in prompt design, managing complex reasoning tasks, and performance evaluation across various scenarios.
+Although LLMs show promising emergent abilities, they frequently lack predictability and control, and their resistance to misleading prompts is still an issue.
+2.2. Prompt Optimization and Selection
 The challenge of optimizing prompts for LLMs has been addressed in several key studies, each contributing unique methodologies to enhance model performance and efficiency.
@@
 (2024) propose a structured three-step methodology that contains data clustering, prompt generation, and evaluation, effectively balancing generality and specificity in prompt selection.
-ProTeGi (Pryzant et al.
-2023) offers a non-parametric approach inspired by gradient descent, leveraging natural language ”gradients” to iteratively refine prompts.
+ProTeGi (Pryzant et al., 2023) offers a non-parametric approach inspired by gradient descent, leveraging natural language "gradients" to iteratively refine prompts.
 Wang et al.
@@
 Zhou et al.
-(2023b) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
-These studies aim to automate prompt design, moving away from traditional manual trial-and-error methods while emphasizing efficiency and scalability across various tasks and models.
-They report significant improvements in LLMs performance, with enhancements ranging from 5% to 31% across different benchmarks.
-This body of work underscores the increasing importance of prompt optimization and selection in unlocking the potential of LLMs and points toward future research avenues, such as exploring theoretical foundations, integrating multiple optimization techniques, and distinguishing between task-specific and general-purpose strategies.
-2.3 Evaluation Benchmarks
+(2023a) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
+These studies seek to automate prompt design, reducing reliance on manual trial-and-error while improving efficiency and scalability across tasks and models.
+They report performance gains of 5% to 31% across benchmarks, highlighting the growing significance of prompt optimization.
+Future research directions include exploring theoretical foundations, combining optimization techniques, and differentiating task-specific from general-purpose strategies.
+2.3. Evaluation Benchmarks
 To facilitate the evaluation and understanding of LLM capabilities, Zhu et al.
@@
 (2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning.
-(Wang et al.
-2021) introduced a multi-task Benchmark for robustness Evaluation of LLMs extends the original GLUE (Wang et al.
-2018) benchmark to assess model robustness against adversarial inputs.
+(Wang et al., 2021) introduced a multitask benchmark for LLM robustness evaluation, which extends the original GLUE (Wang et al., 2018) benchmark to assess model robustness against adversarial inputs.
 It incorporates perturbed versions of existing GLUE tasks, such as paraphrasing, negation, and noise, to test models’ abilities with challenging data.
 The study highlights that despite their success on clean datasets, state-of-the-art models often struggle with adversarial examples, underscoring the importance of robustness evaluations in model development.
-3 Hierarchical Prompting Taxonomy
-Figure 3: Hierarchical Prompting Taxonomy: A taxonomy designed to assess the complexity of prompting strategies based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
-3.1 Governing Rules
+3. Hierarchical Prompting Taxonomy
+Figure 3.
+Hierarchical Prompting Taxonomy: A taxonomy designed to assess the complexity of prompting strategies based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
+3.1. Governing Rules
 Figure 3 illustrates the HPT, a taxonomy that systematically reflects human cognitive functions as outlined in Bloom (1956).
 Each rule embodies complex cognitive processes based on established principles from learning and psychology.
-  1. 1.
-    Basic Recall and Reproduction: This reflects the fundamental cognitive process of remembering and reproducing factual information without analysis or interpretation, which involves mere recognition or retrieval of knowledge from memory (Anderson et al.
-2014).
-  2. 2.
-    Understanding and Interpretation: This corresponds to the second cognitive rule of (Bloom 1956), where individuals must not only recall information but also explain it in their own words, summarize key points or clarify the meaning of content.
+  1. (1)
+    Basic Recall and Reproduction: This reflects the fundamental cognitive process of remembering and reproducing factual information without analysis or interpretation, which involves mere recognition or retrieval of knowledge from memory (Anderson et al., 2014).
+  2. (2)
+    Understanding and Interpretation: This corresponds to the second cognitive rule of (Bloom, 1956), where individuals must not only recall information but also explain it in their own words, summarize key points or clarify the meaning of content.
 This rule demands an intermediate cognitive load involving information processing rather than retrieving it.
-  3. 3.
-    Analysis and Reasoning: This aligns with the analysis stage of (Bloom 1956), which involves higher cognitive functions such as comparison, contrast, and deep understanding of the underlying principles.
+  3. (3)
+    Analysis and Reasoning: This aligns with the analysis stage of (Bloom, 1956), which involves higher cognitive functions such as comparison, contrast, and deep understanding of the underlying principles.
 It is more complex than mere understanding because it requires examining structure and identifying patterns and connections.
-  4. 4.
-    Application of Knowledge and Execution: This mirrors the application and evaluation stages of (Bloom 1956), where individuals must not only understand and analyze but also use knowledge to perform multi-step tasks, solve complex problems, and execute decisions.
+  4. (4)
+    Application of Knowledge and Execution: This mirrors the application and evaluation stages of (Bloom, 1956), where individuals must not only understand and analyze but also use knowledge to perform multi-step tasks, solve complex problems, and execute decisions.
 It represents the most cognitively complex tasks, which require synthesis of information and practical decision-making, highlighting the critical leap from understanding theory to executing it in practice.
 In HPT, the progression from basic recall to application of knowledge reflects increasing cognitive complexity, consistent with educational and cognitive frameworks, where more advanced cognitive processes build on foundational ones, demanding deeper engagement and mental effort.
-3.2 Hierarchical Prompting Framework
+3.2. Hierarchical Prompting Framework
 The HPF consists of five prompting strategies, each assigned a complexity level.
@@
 Section 4.5 examines the hierarchical structure of the HPF in conjunction with the LLM-as-a-Judge framework, validating that the cognitive demands on LLMs can be aligned with those of humans.
-The set of five prompting strategies were chosen from a diverse range of existing strategies to populate the framework, guided by a human judgment policy, prioritizing comprehensiveness in cognitive demands rather than the sheer number of strategies.
-See Appendix A for more details.
-Consequently, the HPF can be replicated or expanded with other relevant prompting strategies that exhibit similar cognitive demands, making the framework adaptable.
-The five prompting strategies, listed from least to most complex, are as follows:
-  1. 1.
-    Role Prompting (Kong et al.
-2024): Prompts that specify the role of the LLM in task resolution represent the lowest level of complexity within the framework, as they exhibit minimal influence from the four principles of the HPT.
-  2. 2.
-    Zero-Shot Chain-of-Thought Prompting (Zero-CoT) (Kojima et al.
-2022): Prompts that use the phrase “Let’s think step by step” without providing prior examples aim to encourage critical thinking and problem-solving.
-This strategy implicitly activates chain-of-thought reasoning, demonstrating moderate influence from rule 3 of the HPT, while exerting a weaker influence from the other rules.
-  3. 3.
-    Three-Shot Chain-of-Thought Prompting (3-CoT) (Wei et al.
-2022b): Prompts that present three examples to guide the LLM’s reasoning process exert a strong influence from rules 1 and 2, requiring recall and interpretation of the examples to complete the task.
-This strategy also demonstrates moderate influence from rule 3, increasing its cognitive demands.
-  4. 4.
-    Least-to-Most Prompting (Zhou et al.
-2023a): A sequence of prompts that breaks the task into sub-problems exerts significant influence from all three rules 1, 2, and 3.
-This approach requires recalling prior prompts, interpreting previous responses, and analyzing them to effectively solve the task, resulting in a highly cognitively demanding strategy.
-  5. 5.
-    Generated Knowledge Prompting (GKP) (Liu et al.
-2022): Prompts that require integrating external knowledge to generate relevant information represent the most complex and cognitively demanding strategy.
-This approach is strongly influenced by rules 2, 3, and 4, as it involves correlating knowledge with the prompt and applying and analyzing external information, making it the most cognitively demanding within the HPT framework.
-In the experiments, Llama-3 8B is used to generate external knowledge.
-3.3 Hierarchical Prompting Index
+The five prompting strategies were selected to ensure comprehensive coverage of cognitive demands rather than maximizing the number of strategies (see Appendix A).
+This makes HPF adaptable, allowing for replication or expansion with similar strategies.
+The strategies, ordered by increasing complexity, are:
+  1. (1)
+    Role Prompting (Kong et al., 2024): Specifies the LLM’s role in task resolution, exerting minimal influence from HPT principles.
+  2. (2)
+    Zero-Shot Chain-of-Thought Prompting (Zero-CoT) (Kojima et al., 2022): Uses “Let’s think step by step” to encourage reasoning, moderately influenced by rule 3.
+  3. (3)
+    Three-Shot Chain-of-Thought Prompting (3-CoT) (Wei et al., 2022b): Provides three examples to guide reasoning, strongly influenced by rules 1 and 2, with moderate influence from rule 3.
+  4. (4)
+    Least-to-Most Prompting (Zhou et al., 2023b): Breaks tasks into sub-problems, requiring recall, interpretation, and analysis, exerting strong influence from rules 1, 2, & 3.
+  5. (5)
+    Generated Knowledge Prompting (GKP) (Liu et al., 2022): Integrates external knowledge, demanding correlation, application, and analysis, making it the most cognitively complex (rules 2, 3, and 4).
+Llama-3 8B generates the external knowledge in experiments.
+3.3. Hierarchical Prompting Index
 HPI is an evaluation metric for assessing the task complexity of LLMs over different datasets, which is influenced by the HPT rules.
-A lower HPI for a dataset suggests that the corresponding LLM is more adept at solving the task with fewer cognition processes.
+A lower HPI for a dataset suggests that the corresponding LLM is more adept at solving the task with fewer cognitive processes.
 For each dataset instance, we begin with the least complex prompting strategy and progressively move through the HPF prompting strategies until the instance is resolved.
 The HPI corresponds to the complexity level of the prompting strategy where the LLM first tackles the instance.
-Algorithm 1 HPI Metric
+Algorithm 1 HPI Computation
 […]
@@
 $\texttt{HPI}=\frac{1}{n}\sum_{j=1}^{n}\texttt{HPI\_List}[j]$
-$m$ is the total number of levels in the HPF, and $n$ is the total number of samples in the evaluation dataset.
-$\texttt{HPI}_{Dataset}$ represents the penalty introduced into the framework by human assessments.
-For further details on human annotation, see Appendix A.
-4 Results
-4.1 Experimental Setup
+Algorithm 1 illustrates the process for determining HPI, with $m$ indicating the total levels within the HPF and $n$ representing the number of samples in the evaluation dataset.
+$\texttt{HPI}_{Dataset}$ denotes the penalty that human evaluations impose on the framework.
+Additional information regarding human annotation is provided in Appendix A.
+4. Results
+4.1. Experimental Setup
 Datasets
-The experiments utilized a diverse set of datasets, including MMLU, GSM8k, HumanEval, BoolQ, CSQA, SamSum, and IWSLT en-fr covering areas such as reasoning, coding, mathematics, question-answering, summarization, and machine translation, to evaluate the framework’s robustness and applicability.
-For further details on evaluation dataset sizes, see Appendix A.
-Reasoning: MMLU (Hendrycks et al.
-2021) includes multiple-choice questions across 57 subjects, covering areas like humanities, social sciences, physical sciences, basic mathematics, U.S.
-history, computer science, and law.
-CommonSenseQA (CSQA) (Talmor et al.
-2019) contains  12,000 questions to assess commonsense reasoning.
-Coding: HumanEval (Chen et al. 2021a) features 164 coding challenges, each with a function signature, docstring, body, and unit tests, designed to avoid training data overlap with LLMs.
-Mathematics: Grade School Math 8K (GSM8k) (Cobbe et al. 2021) comprises 8.5K diverse math problems for multi-step reasoning, focusing on basic arithmetic and early Algebra.
-Question-Answering: BoolQ (Clark et al. 2019) consists of  16,000 True/False questions based on Wikipedia passages for binary reading comprehension.
-Summarization: SamSum (Gliwa et al. 2019) features  16,000 human-generated chat logs with summaries for dialogue summarization.
-Machine Translation: IWSLT-2017 en-fr (IWSLT) (Cettolo et al. 2017) is a parallel corpus with thousands of English-French sentence pairs from TED Talks for translation tasks.
-Large Language Models
-For the evaluation, LLMs with parameter sizes ranging from 7 billion to 12 billion from top open-source models and top proprietary models were selected to determine the effectiveness of the proposed framework across varied parameter scales and architectures.
-Proprietary Models: GPT-4o (OpenAI 2024), Claude 3.5 Sonnet (Anthropic 2024)
-Open-Source models: Gemma 7B (Team et al. 2024a), Mistral 7B (Jiang et al. 2023), Llama-3 8B (AI@Meta 2024), Gemma-2 9B (Team et al. 2024b) and Mistral-Nemo 12B (Mistral AI and NVIDIA 2024).
+We evaluated the framework on diverse datasets spanning reasoning, coding, mathematics, question-answering, summarization, and machine translation. For dataset sizes, see Appendix A.
+Reasoning: MMLU (Hendrycks et al., 2021) (57 subjects, multiple-choice), CSQA (Talmor et al., 2019) ( 12K commonsense questions).
+Coding: HumanEval (Chen et al., 2021a) (164 function-based coding tasks).
+Mathematics: GSM8k (Cobbe et al., 2021) (8.5K multi-step math problems).
+Question-Answering: BoolQ (Clark et al., 2019) ( 16K True/False questions from Wikipedia).
+Summarization: SamSum (Gliwa et al., 2019) ( 16K human-annotated dialogue summaries).
+Machine Translation: IWSLT-2017 en-fr (Cettolo et al., 2017) (TED Talk parallel corpus).
+Large Language Models: We tested LLMs ranging from 7B to 12B parameters across open-source and proprietary models.
+Proprietary LLMs: GPT-4o (OpenAI, 2024), Claude 3.5 Sonnet (Anthropic, 2024).
+SLMs: Gemma 7B (Team et al., 2024a), Mistral 7B (Jiang et al., 2023), Llama-3 8B (AI@Meta, 2024), Gemma-2 9B (Team et al., 2024b), Mistral-Nemo 12B (Mistral AI and NVIDIA, 2024).
 Additional Evaluation Metrics
-  • Coding: The Pass@k (Chen et al. 2021b) metric measures the probability of at least one correct solution among the top k outputs, used for evaluating code generation.
-  • Summarization: ROUGE-L (Lin 2004) evaluates the longest common subsequence between generated text and reference, focusing on sequence-level similarity for summaries and translations.
-  • Machine Translation: BLEU (Papineni et al. 2002) is a precision-based metric that assesses machine-generated text quality by comparing n-grams with reference texts.
-In the experiments, thresholds of 0.15, and 0.20 were established for summarization and machine translation tasks to define the conditions required for task completion at each complexity level of the HPF.
-These thresholds allowed for iterative refinement of HPF prompting strategies.
-4.2 Results on Standard Benchmarks: MMLU, GSM8K, and Humaneval
+Coding
+: Pass@k (Chen et al., 2021b) estimates the probability of at least one correct solution among the top k outputs for code generation.
+Summarization: ROUGE-L (Lin, 2004) measures sequence-level similarity via the longest common subsequence.
+Machine Translation: BLEU (Papineni et al., 2002) evaluates n-gram precision against reference texts.
+Summarization and translation tasks used thresholds of 0.15 and 0.20, respectively, to define task completion at each HPF complexity level, enabling iterative refinement of prompting strategies.
+Figure 4.
+Performance Comparison of HPT-based Evaluation vs.
+Standard Evaluation: Performance improvements (in %) when using HPT-based evaluation compared to standard evaluation across three benchmarks: MMLU, GSM8k, and HumanEval.
+Positive values indicate performance gains with HPT, while negative values indicate performance decreases.
+The baseline standard evaluation scores are sourced from Hugging Face leaderboard and official research reports.
+4.2. Results on Standard Benchmarks: MMLU, GSM8K, and Humaneval
 The evaluation of HPF effectiveness as shown in Figure 4 spans three standard benchmarks: MMLU, GSM8k, and HumanEval.
@@
 Mistral 7B achieved an exception 62.5% improvement in HumanEval scores, followed by Mistral-Nemo 12B with an impressive 51.4% improvement, and Gemma-2 9B with a 50.8% enhancement.
-These findings indicate that HPF improves performance across all benchmarks for most of the LLMs, its impact is particularly pronounced in programming tasks, suggesting that the technique may be especially valuable for enhancing code-related capabilities.
-Table 1 highlights the improved performance of various LLMs on MMLU, with all models showing an HPI index below three.
-This indicates that reasoning over most MMLU samples requires minimal cognitive effort for these models, compared to baseline multi-shot CoT methods (5 shot), which typically require more than five examples and are more cognitively demanding according to HPT.
-Interestingly, while Claude 3.5 Sonnet achieves the highest MMLU accuracy, GPT-4o records the best HPI score, showing that minimal cognitive effort does not necessarily equate to the best performance.
-The enhancement in GSM8k is relatively smaller compared to MMLU, with decreased performances for both Mistral 7B and Gemma 7B.
-The high HPI values for Gemma 7B and Mistral 7B indicate that none of the five prompting strategies in HPF posed significant cognitive challenges for these LLMs, highlighting a limitation of the HPF.
-As shown in Table 2, Claude 3.5 Sonnet achieves a perfect pass@1 of 1.00 with low HPI values, outperforming GPT-4o, which scores 0.95 but has a higher HPI.
-Gemma 7B struggles with the lowest pass@1 of 0.79 and the highest HPI of 3.71, indicating a need for more complex prompting strategy.
-Interestingly, HPF significantly enhanced the performance of most LLMs across three benchmark datasets, even when the HPI difference was less than 1 relative to the best performing LLMs.
-This highlights that tailoring the prompting strategy to align with the complexity of each dataset instance can lead to substantial improvements, achieving performance levels comparable to state-of-the-art LLMs such as GPT-4o and Claude 3.5 Sonnet on these benchmarks.
-Figure 4: Performance Comparison of HPT-based Evaluation vs.
-Standard Evaluation: Performance improvements (in %) when using HPT-based evaluation compared to standard evaluation across three benchmarks: MMLU, GSM8k, and HumanEval.
-Positive values indicate performance gains with HPT, while negative values indicate performance decreases.
-The baseline standard evaluation scores are sourced from Hugging Face leaderboard and official research reports.
+The results suggest that HPF enhances performance on all benchmarks for the majority of SLMs and achieves similar performance to LLMs such as GPT-4o and Claude 3.5 Sonnet, thereby addressing RQ1, its impact is particularly pronounced in programming tasks, suggesting that the technique may be especially valuable for enhancing code-related capabilities.
+Table 1.
+HPI (lower is better) and accuracy of LLMs across MMLU, GSM8K, BoolQ, and CSQA datasets.
+Blue indicates datasets where the LLM with the best HPI does not achieve the best performance.
+Green indicates the LLM with the best performance over the maximum number of datasets.
 DATASETS MMLU GSM8k BoolQ CSQA
@@
 Mistral 7B 2.89 81.45 5.11 46.93 1.41 98.07 2.49 82.06
-Table 1: HPI (lower is better) and accuracy of LLMs across MMLU, GSM8K, BoolQ, and CSQA datasets.
+Table 1 highlights the improved performance of various LLMs on MMLU, with all models showing an HPI index below three.
+This indicates that reasoning over most MMLU samples requires minimal cognitive effort for these models, compared to baseline multi-shot CoT methods (5 shot), which typically require more than five examples and are more cognitively demanding according to HPT.
+Interestingly, while Claude 3.5 Sonnet achieves the highest MMLU accuracy, GPT-4o records the best HPI score, showing that minimal cognitive effort does not necessarily equate to the best performance addressing RQ2.
+The enhancement in GSM8k is relatively smaller compared to MMLU, with decreased performances for both Mistral 7B and Gemma 7B.
+The high HPI values for Gemma 7B and Mistral 7B indicate that none of the five prompting strategies in HPF posed significant cognitive challenges for these LLMs, i.e more cognitively demanding prompting strategies are needed, highlighting a limitation of the HPF.
+As shown in Table 2, Claude 3.5 Sonnet achieves a perfect pass@1 of 1.00 with low HPI values, outperforming GPT-4o, which scores 0.95 but has a higher HPI.
+Gemma 7B struggles with the lowest pass@1 of 0.79 and the highest HPI of 3.71, indicating a need for a more complex prompting strategy.
+Notably, HPF noticeably boosted the performance of the majority of LLMs on three benchmark datasets, despite the HPI difference being less than 1 compared to the top-performing LLMs.
+This suggests that even with a minimal number of inferences, utilizing HPF can achieve optimal performance, unlike multi-shot prompting and prompt optimization strategies, thereby addressing RQ3.
+This highlights that tailoring the prompting strategy to align with the complexity of each dataset instance can lead to substantial improvements, achieving performance levels comparable to state-of-the-art LLMs such as GPT-4o and Claude 3.5 Sonnet on these benchmarks.
+Table 2.
+HPI (lower is better) and Pass@1 of LLMs on the HumanEval dataset.
 Blue indicates datasets where the LLM with the best HPI does not achieve the best performance.
-Green indicates the LLM with the best performance over the maximum number of datasets.
+Green indicates the LLMs with the best performance over the dataset.
 DATASET HumanEval
@@
 Mistral 7B 1.10 0.93
-Table 2: HPI (lower is better) and Pass@1 of LLMs on the HumanEval dataset.
-Blue indicates datasets where the LLM with the best HPI does not achieve the best performance.
-Green indicates the LLMs with the best performance over the dataset.
-Figure 5: Hierarchy of prompting strategies with LLM-as-a-Judge framework with GPT-4o as the judge.
-4.3 Results on Other Datasets
-Table 1 presents the performance of LLMs on the BoolQ and CSQA datasets.
-Notably, no significant insights emerge from the results, aside from GPT-4o performing unexpectedly poorly, which contrasts with its typical performance.
-With most LLMs achieving near-perfect scores, the BoolQ dataset appears to lack the complexity needed to serve as an effective benchmark for modern LLMs, as they perform exceptionally well even with minimal cognitive prompting strategies.
-This underscores the utility of HPF in evaluating dataset complexities relative to an LLM, offering researchers valuable insights for designing more challenging and robust benchmarks.
-Table 3 presents the performance of LLMs on IWSLT and SamSum datasets at varying thresholds.
-GPT-4o consistently achieved the highest scores across all thresholds, while most models, except Gemma 7B, performed similarly.
-Interestingly, Claude 3.5 Sonnet, which excelled in reasoning tasks, did not perform as strongly in summarization and translation tasks.
-The threshold selection is guided by the observed performance plateau across most LLMs as the threshold increases.
-For a detailed explanation of the threshold selection process, please refer to Appendix B.
+4.3. Results on Other Datasets
+Table 1 presents LLM performance on the BoolQ and CSQA datasets.
+While no significant insights emerge, an unexpected result is GPT-4o’s poor performance, which deviates from its typical trend.
+With most LLMs achieving near-perfect scores, BoolQ appears insufficiently complex to serve as an effective benchmark for modern LLMs, as they excel even with minimal cognitive prompting.
+This highlights HPF’s value in assessing dataset complexity relative to LLM capabilities, providing researchers with insights for designing more challenging and robust benchmarks.
+Table 3. HPI (lower is better), BLEU score for IWSLT, and ROUGE-L score for SamSum, of LLMs with thresholds.
 DATASETS IWSLT SamSum
@@
 Mistral 7B 3.52 4.14 0.22 0.22 1.26 1.68 0.21 0.22
-Table 3: HPI (lower is better), BLEU score for IWSLT, and ROUGE-L score for SamSum, of LLMs with thresholds.
-4.4 Complexity Levels with LLM-as-a-Judge
+Table 3 presents the performance of LLMs on IWSLT and SamSum datasets at varying thresholds.
+GPT-4o consistently achieved the highest scores across all thresholds, while most models, except Gemma 7B, performed similarly.
+Interestingly, Claude 3.5 Sonnet, which excelled in reasoning tasks, did not perform as strongly in summarization and translation tasks.
+The threshold selection is guided by the observed performance plateau across most LLMs as the threshold increases.
+4.4. Threshold Selection for SamSum and IWSLT
+In addition to the 0.15 and 0.20 thresholds presented in the main experiments, extended evaluations were conducted on the IWSLT and SamSum datasets using thresholds of 0.25 and 0.30 with GPT-4o, Mistral-Nemo 12B, and Llama-3 to assess the impact of varying thresholds on LLM performance.
+SamSum Dataset: In the summarization task, increasing the threshold evaluates an LLM’s ability to condense content while retaining key information.
+Higher thresholds like 0.25 and 0.30 reveal the trade-offs between conciseness and informativeness.
+However, as shown in Figure 5, there was no significant improvement in ROUGE-L, except for a slight increase with GPT-4o.
+The experiments showed a sharp rise in HPI, reflecting the increased task complexity.
+These results suggest that LLM performance has plateaued, with no further gains at higher thresholds.
+This validates that the use of 0.15 and 0.20 thresholds are sufficient for optimal LLM performance.
+Figure 5. Comparison of HPI and ROUGE-L scores across different threshold values on SamSum dataset.
+IWSLT Dataset: In machine translation, higher thresholds (0.25 and 0.30) impose stricter evaluations, assessing how well models capture the nuances of the source text.
+Lower thresholds (0.15 and 0.20) focus on general adequacy, while higher ones test performance under more challenging conditions.
+As shown in Figure 6, no BLEU improvements were observed across any LLMs, with models either reaching saturation or showing decreased performance alongside a rapid rise in HPI.
+This validates the selection of 0.15 and 0.20 thresholds are sufficient for optimal LLM performance.
+Figure 6. Comparison of HPI and BLEU score across different threshold values in the translation task.
+4.5. Complexity Levels with LLM-as-a-Judge
 This study evaluated prompting strategies by assessing how GPT-4o, as the LLM judge, replicates the hierarchical complexity levels of these strategies using a systematic scoring approach across tasks.
@@
 Figure 8 shows the scoring distribution across the four HPT rules for each strategy.
-Further details related to dataset specifications and scoring method are in Appendix C.
-Figure 6: Scoring distribution for each of the four rules of the HPT for the prompting strategies in the HPF.
-4.5 Parallels with System 1 and System 2 Thinking
-HPF align closely with the principles of System 1 and System 2 thinking from dual-process cognitive theories (Booch et al.
-2021).
-HPT categorizes tasks and HPF structures prompts based on their cognitive complexity, mirroring how humans allocate cognitive resources.
-For tasks with low cognitive demands, HPF employs simple prompts that parallel System 1 thinking.
-These tasks, like fact recall or basic classification, require minimal reasoning, allowing the LLM to respond quickly and efficiently without extensive computation.
-For instance, asking an LLM to ”identify the capital of a country” is analogous to a person retrieving a familiar fact using System 1.
-In contrast, tasks with high cognitive demands involve prompts that guide the LLM through complex reasoning, abstraction, or multi-step problem-solving—analogous to System 2 thinking.
-Examples include generating logical arguments or solving intricate problems, where deliberate and resource-intensive processes are necessary.
-Just as System 2 engages when a problem exceeds the capacity of System 1, higher levels of HPF are invoked for tasks requiring deeper analysis.
+Further details related to evaluation dataset specifications and scoring method are in Appendix B.
+Figure 7. Hierarchy of prompting strategies with LLM-as-a-Judge framework with GPT-4o as the judge.
+Figure 8. Scoring distribution for each of the four rules of the HPT for the prompting strategies in the HPF.
+4.6. Parallels with System 1 and System 2 Thinking
+HPF parallels dual-process cognitive theories’ System 1 and System 2 thinking (Booch et al., 2021; Kahneman, 2011).
+HPT classifies tasks, and HPF designs prompts based on cognitive complexity, reflecting human cognitive resource allocation.
+For tasks with low cognitive demands, HPF uses simple prompts akin to System 1 thinking, like fact recall or basic classification, enabling quick LLM responses with minimal reasoning.
+Conversely, tasks with high cognitive demands require prompts for complex reasoning and problem-solving, similar to System 2 thinking, involving logical arguments or intricate problems needing deliberate processing.
+Elevated HPF levels are used for tasks demanding deep analysis.
 HPF explicitly measures this transition with HPI, assessing the cognitive load required for each task.
@@
 This parallel highlights how HPT bridges computational strategies with human-like cognitive models, enabling more nuanced task evaluation and resource allocation.
-4.6 Adaptive HPF
+4.7. Adaptive HPF
 The Adaptive HPF automates the selection of the optimal complexity level in the HPF using a prompt-selector, Llama-3 8B in a zero-shot setting, bypassing iterative steps.
@@
 This suggests that Adaptive HPF struggles to select the optimal complexity level, likely due to hallucinations by the prompt-selector when choosing the prompting strategy.
-For more results and ablation studies, see Appendix E.
-Figure 7: HPI of datasets for LLMs in Adaptive HPF.
-5 Conclusion
-The HPT provides a strong and efficient approach for assessing LLMs by focusing solely on the cognitive demands of different tasks.
-The results emphasize that the HPF is effective in evaluating diverse datasets, using the most cognitively effective prompting strategies tailored to task complexity, which results in improved LLM performance across multiple datasets.
-This method not only offers an in-depth insight into LLM’s problem-solving abilities but also suggests that dynamically choosing suitable prompting strategies can enhance LLM performance, setting the stage for future improvements in LLM evaluation methods.
-This study paves the way for formulating and designing evaluation methods grounded in human cognitive principles, aligning them with the problem-solving capabilities of LLMs.
-Additionally, it facilitates the development of more robust benchmarks and in-context learning methodologies to effectively assess LLM performance across various tasks.
-6 Ethical Statement
-The $\texttt{HPI}_{Dataset}$ assigned by experts to the datasets: MMLU, GSM8k, Humaneval, BoolQ, CSQA, IWSLT, and SamSum may introduce bias into the comparative analysis.
-This potential bias stems from the subjective nature of expert scoring, which can be influenced by individual experience and perspective.
-Despite this, the datasets utilized in this study are publicly available and widely recognized in the research community, thereby minimizing the risk of unanticipated ethical issues.
-Nevertheless, it is crucial to acknowledge the possibility of scoring bias to ensure transparency and integrity in the analysis.
+For more results and ablation studies, see Appendix C.
+Figure 9. HPI of datasets for LLMs in Adaptive HPF.
+The prompt-selector can dynamically select the most suitable prompting strategy for a given task’s complexity from the HPF’s hierarchy of complexity levels.
+To determine the most effective prompting strategy to complete the task, the prompt-selector was given a maximum number of iterations equivalent to the number of levels in the manual HPF.
+The score for ith iteration is i + x, where x is the complexity level by the prompt-selector.
+If the LLM fails to complete the task after all iterations, it is assigned a penalty, $\texttt{HPI}_{Dataset}$.
+Algorithm 2 HPI Computation for Adaptive HPF
+[…]
+for sample $j$ in the evaluation dataset do
+[…]
+     for iteration $i=1$ to $m$ do
+         Select prompting strategy at level $x$
+         if LLM completes the task at iteration $i$ then
+[…]
+[…]
+              break
+         end if
+     end for
+     if solved = False then
+[…]
+     end if
+end for
+$\texttt{HPI}_{\texttt{Adaptive}}=\frac{1}{n}\sum_{j=1}^{n}\texttt{HPI\_List}[j]$
+Algorithm 2 demonstrates the calculation of HPI for an adaptive HPF, where $x$ denotes the HPF level chosen by the prompt-selector at the $i$th iteration as the task is being tackled.
+Here, $m$ indicates the total number of HPF levels, and $n$ signifies the total quantity of samples in the evaluation set.
+5. Conclusion
+The HPT offers an efficient way to evaluate LLMs by focusing on task cognitive demands.
+It shows that cognitively inspired selection of prompting strategies enhances LLM performance across various datasets.
+This method offers insights into LLM problem-solving and improves evaluation methods based on human cognition, supporting better in-context learning strategies for assessing LLMs.
+6. Limitations
+Human Annotation Constraints: A limitation of this study is the reliance on human evaluation for inducing the $\texttt{HPI}_{Dataset}$ penalty into the HPF.
+While this study assessed 5% of the datasets, expanding coverage would offer a more comprehensive analysis.
+However, due to constraints in human resources for manual annotation, we could not include a larger portion.
+Future work could address this by increasing manpower or automating parts of the evaluation process.
+HPF Optimization: The effectiveness of the HPF heavily relies on the quality of the prompts used at each level of the taxonomy.
+Crafting high-quality prompts that accurately reflect the subtleties of each level demands considerable expertise and repeated refinement.
+This study only investigated a limited set of prompting strategies within the HPF, indicating a need for further research into creating diverse structural frameworks and incorporating additional prompting strategies.
+Zero-shot Prompt Selection: HPF dynamically determines the optimal cognitive complexity level by iterating through the framework’s levels, which leads to increased inference time.
+While this study investigated Adaptive HPF for zero-shot prompt selection, it faced considerable hallucinations.
+Future research should focus on automating HPF using fine-tuning or reinforcement learning-based approaches to select the appropriate complexity level without manual iteration.
+This strategy would optimize inference time and improve overall performance.
+7. Ethical Statement
+The $\texttt{HPI}_{Dataset}$ assigned by experts to MMLU, GSM8k, HumanEval, BoolQ, CSQA, IWSLT, and SamSum may introduce bias due to the subjective nature of expert scoring, influenced by individual experience and perspective.
+However, these publicly available, widely recognized datasets help mitigate unforeseen ethical concerns.
+Acknowledging potential scoring bias remains essential for transparency and integrity in the analysis.
 Appendix A Human Annotation and Judgement Policy
-A.1 Human Annotation Policy
+A.1. Human Annotation Policy
 $\texttt{HPI}_{Dataset}$ is introduced to penalize the HPI of tasks or samples unsolvable by the LLM, aligning the framework more closely with human cognitive demands and enhancing its comprehensiveness.
@@
 Each sample was scored on a scale from 1 (lowest complexity level) to 5 (highest complexity level) for the following criteria:
-  1. 1.
+  1. (1)
     Basic Understanding and Reproduction: This criterion evaluates the annotator’s ability to comprehend and accurately reproduce the content.
-  2. 2.
+  2. (2)
     Understanding and Interpretation: This criterion assesses the annotator’s depth of understanding and the ability to interpret the information correctly.
-  3. 3.
+  3. (3)
     Analysis and Reasoning: This criterion measures the annotator’s ability to analyze the information and apply logical reasoning.
-  4. 4.
+  4. (4)
     Application of Knowledge and Execution: This criterion evaluates the annotator’s practical application of knowledge and the execution of tasks based on the relevant knowledge.
@@
 Selecting 5% of the dataset as the representative set size balances quality assessment and feasibility, capturing the dataset’s diversity and ensuring that human annotations encompass a broad range of cases without needing to annotate every sample.
+Table 4.
+$\texttt{HPI}_{Dataset}$ scores across datasets evaluated by human annotators.
+The table lists the evaluation set size, representative set size, and $\texttt{HPI}_{Dataset}$ for various datasets.
+$\texttt{HPI}_{Dataset}$ scores provide a measure of task complexity relative to human annotators.
 Dataset Evaluation Set Size Representative Set Size $\texttt{HPI}_{Dataset}$
@@
 SamSum 819 40 2.23
-Table 4: $\texttt{HPI}_{Dataset}$ scores across datasets evaluated by human annotators.
-The table lists the evaluation set size, representative set size, and $\texttt{HPI}_{Dataset}$ for various datasets.
-$\texttt{HPI}_{Dataset}$ scores provide a measure of task complexity relative to human annotators.
-A.2 Human Judgement Policy
+A.2. Human Judgement Policy
 To populate the HPF with relevant prompting strategies across a wide range of strategies, human annotators who adhered to the annotation policy for assessing $\texttt{HPI}_{Dataset}$ were instructed to follow a judgment policy for a predefined set of prompting strategies.
@@
 GKP L H H H
-Table 5: Human judgment of influence of the rules of taxonomy on different prompting strategies in solving the tasks of the representative set.
+Table 5.
+Human judgment of influence of the rules of taxonomy on different prompting strategies in solving the tasks of the representative set.
 The ratings are provided based on a voting system involving all human annotators.
 Green represents the prompting strategies selected for populating the complexity levels of the HPF.
-Appendix B Threshold Selection for Summarization and Translation Tasks
-In addition to the 0.15 and 0.20 thresholds presented in the main paper, extended evaluations were conducted on the IWSLT and SamSum datasets using thresholds of 0.25 and 0.30 with GPT-4o, Mistral-Nemo 12B, and Llama-3 to assess the impact of varying thresholds on LLM performance.
-B.1 Summarization Task
-In the summarization task, increasing the threshold evaluates an LLM’s ability to condense content while retaining key information.
-Higher thresholds like 0.25 and 0.30 reveal the trade-offs between conciseness and informativeness.
-However, as shown in Figure 8, there was no significant improvement in ROUGE-L, except for a slight increase with GPT-4o.
-The experiments showed a sharp rise in HPI, reflecting the increased task complexity.
-These results suggest that LLM performance has plateaued, with no further gains at higher thresholds.
-This validates the use of 0.15 and 0.20 thresholds in the main paper are sufficient for optimal LLM performance.
-Figure 8: Comparison of HPI and ROUGE-L score across different threshold values in summarization task.
-B.2 Translation Task
-In machine translation, higher thresholds (0.25 and 0.30) impose stricter evaluations, assessing how well models capture the nuances of the source text.
-Lower thresholds (0.15 and 0.20) focus on general adequacy, while higher ones test performance under more challenging conditions.
-As shown in Figure 9, no BLEU improvements were observed across any LLMs, with models either reaching saturation or showing decreased performance alongside a rapid rise in HPI.
-This validates the selection of 0.15 and 0.20 thresholds in the main paper as sufficient for optimal LLM performance.
-Figure 9: Comparison of HPI and BLEU score across different threshold values in the translation task.
-Appendix C LLM-as-a-Judge
-C.1 Scoring Prompt Template
+Appendix B LLM-as-a-Judge
+B.1. Scoring Prompt Template
 The system prompt is designed to guide the LLM judge in evaluating different prompting strategies based on four specific criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Execution.
@@
 Range: 1-5
-C.2 Hybrid Dataset
+B.2. Hybrid Dataset
 The hybrid dataset is composed of 1106 samples uniformly distributed over seven different task-specific datasets, covering a wide range of language understanding and generation tasks.
@@
 The evaluation uses a hybrid dataset composed of samples from various task-specific datasets and each dataset contributes specific types of tasks:
-  1. 1.
+  1. (1)
     MMLU (Massive Multitask Language Understanding)
-  2. 2.
+  2. (2)
     HumanEval (Code Generation and Completion)
-  3. 3.
+  3. (3)
     GSM8K (Grade School Math 8K)
-  4. 4.
+  4. (4)
     BoolQ (Boolean Questions)
-  5. 5.
+  5. (5)
     CSQA (Commonsense Question Answering)
-  6. 6.
+  6. (6)
     IWSLT (International Workshop on Spoken Language Translation)
-  7. 7.
+  7. (7)
     SamSum (Dialogue Summarization)
-C.3 Scoring Method
+B.3. Scoring Method
 For each prompting strategy (Role Prompting, Zero-shot CoT, Three-shot CoT, Least to Most Prompting, Generated Knowledge Prompting), the system:
-  1. 1.
+  1. (1)
     Applies the prompting strategy to each sample in the hybrid dataset
-  2. 2.
+  2. (2)
     Generates an answer using GPT-4o
-  3. 3.
+  3. (3)
     Presents the prompt, generated answer, and correct answer to the LLM judge
-  4. 4.
+  4. (4)
     Collects scores for each of the four criteria and the system calculates average scores for each criterion across all tasks and datasets.
 This study ensured that both the human judge and the LLM judge utilized the same scoring methodology to eliminate any potential bias in the comparison.
-Appendix D Limitations
-Human Annotation Constraints: A limitation of this study is the reliance on human evaluation for inducing the $\texttt{HPI}_{Dataset}$ penalty into the HPF.
-While this study assessed 5% of the datasets, expanding coverage would offer a more comprehensive analysis.
-However, due to constraints in human resources for manual annotation, we could not include a larger portion.
-Future work could address this by increasing manpower or automating parts of the evaluation process.
-HPF Optimization: The effectiveness of the HPF heavily relies on the quality of the prompts used at each level of the taxonomy.
-Crafting high-quality prompts that accurately reflect the subtleties of each level demands considerable expertise and repeated refinement.
-This study only investigated a limited set of prompting strategies within the HPF, indicating a need for further research into creating diverse structural frameworks and incorporating additional prompting strategies.
-Zero-shot Prompt Selection: HPF dynamically determines the optimal cognitive complexity level by iterating through the framework’s levels, which leads to increased inference time.
-While this study investigated Adaptive HPF for zero-shot prompt selection, it faced considerable hallucinations.
-Future research should focus on automating HPF using fine-tuning or reinforcement learning-based approaches to select the appropriate complexity level without manual iteration.
-This strategy would optimize inference time and improve overall performance.
-Appendix E Adaptive HPF
-E.1 HPI for Adaptive HPF
-The prompt-selector can dynamically select the most suitable prompting strategy for a given task’s complexity from the HPF’s hierarchy of complexity levels.
-To determine the most effective prompting strategy to complete the task, the prompt-selector was given a maximum number of iterations equivalent to the number of levels in the manual HPF.
-The score for ith iteration is i + x, where x is the complexity level by the prompt-selector.
-If the LLM fails to complete the task after all iterations, it is assigned a penalty, $\texttt{HPI}_{Dataset}$.
-Algorithm 2 Prompt-Selector for Adaptive HPF
-[…]
-for sample $j$ in the evaluation dataset do
-[…]
-     for iteration $i=1$ to $m$ do
-         Select prompting strategy at level $x$
-         if LLM completes the task at iteration $i$ then
-[…]
-[…]
-              break
-         end if
-     end for
-     if solved = False then
-[…]
-     end if
-end for
-$\texttt{HPI}_{\texttt{Adaptive}}=\frac{1}{n}\sum_{j=1}^{n}\texttt{HPI\_List}[j]$
-$x$ represents the level of the HPF selected by prompt-selector at $i$th iteration at which the task is addressed, $m$ represents the total number of levels in the HPF, and $n$ denotes the total number of samples in the evaluation set.
+Table 6. HPI (lower is better) of LLMs across datasets (with thresholds) for Adaptive HPF.
 Model BoolQ CSQA IWSLT (0.15) IWSLT (0.20) SamSum (0.15) SamSum (0.20)
@@
 Gemma 7B 5.1514 5.5771 6.5947 6.6605 5.7229 6.4347
-Table 6: HPI (lower is better) of LLMs across datasets (with thresholds) for Adaptive HPF.
+Table 7. Performance scores of LLMs across datasetsfor Adaptive HPF.
 Dataset Metric Threshold Llama-3 8B Phi-3 3.8B Mistral 7B Gemma 7B
@@
 0.2 0.24981 0.21580 0.28335 0.16179
-Table 7: Performance scores of LLMs across datasetsfor Adaptive HPF.
-E.2 Hallucination in Adaptive HPF
+Appendix C Hallucination in Adaptive HPF
 Hallucinations in prompt-selector refer to instances where the LLM generates incorrect or misleading prompting levels or nonsensical information that is not supported by the HPF.
 These hallucinations can occur across various tasks, including question answering, multiple-choice questions, translation, and summarization.
-For the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations (represented by ’…’).
+For the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations.
 However, by the fourth iteration, Llama-3 8B manages to answer correctly at Level 2.
@@
 While the LLM can eventually produce correct answers, as seen in the BoolQ and SamSum tasks, it often requires multiple attempts and may still fail in tasks like IWSLT translation.
-E.3 Prompt Template for Prompt-Selector
+C.1. Prompt Template for Prompt-Selector
 The prompt-selector in adaptive HPF selects the prompting level based on the task complexity to address the task.
@@
 Select only the index (do not provide the name) of the most effective prompting strategy.
-Appendix F Computational Budget
+Appendix D Computational Budget
 All evaluation experiments and ablation studies were conducted on V100 GPUs (16GB and 32GB variants), utilizing a total of around 9,000 computation hours, this equates to approximately 1.125 petaflop-hours of computational resources.
-Appendix G Large Language Models Used for Evaluation
+Appendix E Large Language Models Used for Evaluation
 The HPF supports leading open source and proprietary LLMs and includes mechanisms for optimizing performance through advanced quantization techniques.
 The experiments were conducted on the following instruction-tuned LLMs, and the model description and licenses are discussed in Table 8.
+Table 8. License information for LLMs used in the experiments.
 Model License Type Usage Restrictions
@@
 Phi-3 3.8B Open-source Broad use allowed, must include original license and notices
-Table 8: License information for LLMs used in the experiments.
 The LLMs were loaded in 4-bit precision format, with a maximum generation limit of 1024 tokens per run to ensure concise outputs.
@@
 Additionally, a repetition penalty was applied to discourage the generation of repeated phrases, promoting coherent and varied text output.
-Appendix H Prompt Templates
-H.1 Level 1: Role Prompting
+Appendix F Prompt Templates
+F.1. Level 1: Role Prompting
 Role prompting represents the most basic interaction with an LLM, assigning it a specific role or task without additional context or examples.
@@
 Table 9 shows the prompt templates used for role prompting across all datasets in the experiments.
+Table 9. Prompt templates of different datasets for Role Prompting.
 Dataset Prompt
@@
 MMLU Choose the answer: ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4" as a critical thinker.
-Table 9: Prompt templates of different datasets for Role Prompting.
-H.2 Level 2: Zero-shot Chain-of-Thought Prompting
+F.2. Level 2: Zero-shot Chain-of-Thought Prompting
 Zero-shot Chain-of-Thought (CoT) prompting enhances basic role prompting by requiring the LLM to generate a reasoning process for a task, despite not being explicitly trained on similar examples.
@@
 Table 10 displays the prompt templates used for Zero-CoT across all datasets in the experiments.
+Table 10. Prompt templates of different datasets for Zero-shot Chain-of-Thought Prompting.
 Dataset Prompt
@@
 MMLU Choose the answer: ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4". Let’s think step by step.
-Table 10: Prompt templates of different datasets for Zero-shot Chain-of-Thought Prompting.
-H.3 Level 3: Three-Shot Chain-of-Thought Prompting
+F.3. Level 3: Three-Shot Chain-of-Thought Prompting
 Three-shot Chain-of-Thought (CoT) prompting builds on the zero-shot approach by providing the LLM with three task examples, including the reasoning steps used to reach the solution.
@@
 Table 11 shows the prompt templates used for 3-CoT across all datasets in the experiments.
+Table 11. Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting.
 Dataset Prompt
@@
 ‘‘option 4".
-Table 11: Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting.
-H.4 Level 4: Least-to-Most Prompting
+F.4. Level 4: Least-to-Most Prompting
 Least-to-most prompting is an advanced technique that gradually increases prompt complexity, starting with simpler tasks and progressing to more complex challenges.
@@
 Table 12 displays the prompt templates used for Least-to-Most Prompting across all datasets in the experiments.
+Table 12. Prompt templates of different datasets for Least-to-Most Prompting.
 Dataset Prompt
@@
 "option 4".
-Table 12: Prompt templates of different datasets for Least-to-Most Prompting.
-H.5 Level 5: Generated Knowledge Prompting
+F.5. Level 5: Generated Knowledge Prompting
 Generated Knowledge prompting is one of the most complex techniques in HPF, where the LLM not only addresses the task but also integrates relevant additional information to enhance its response.
@@
 Table 13 shows the prompt templates used for Generated Knowledge Prompting across all datasets in the experiments.
+Table 13. Prompt templates of different datasets for Generated Knowledge Prompting.
 Dataset Prompt
@@
 MMLU Choose the answer. "question", options: A. "option 1" B. "option 2" C. "option 3" D. "option 4" using knowledge of the question:"pred"
-Table 13: Prompt templates of different datasets for Generated Knowledge Prompting.